### v1

# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

Abstract
The adoption of electric vehicles (EVs) is considered critical to achieving climate goals, yet it hinges on consumer interest. This study explores how public intent to purchase EVs relates to four unexamined factors: exposure to EV information, perceptions of EVs’ environmental benefits, views on government climate policy, and confidence in future EV infrastructure; while controlling for prior EV ownership, political affiliation, and demographic characteristics (e.g., age, gender, education, and geographic location). We utilize data from three nationally representative opinion polls conducted by the Pew Research Center between 2021 and 2023, and employ Bayesian techniques to estimate the ordinal probit and ordinal quantile models. Results from ordinal probit show that respondents who are well-informed about EVs, perceive them as environmentally beneficial, or are confident in development of charging stations are more likely to express strong interest in buying an EV, with covariate effects--a metric rarely reported in EV research--of 10.2, 15.5, and 19.1 percentage points, respectively. In contrast, those skeptical of government climate initiatives are more likely to express no interest, by more than 10 percentage points. Prior EV ownership exhibits the highest covariate effect (ranging from 19.0 to 23.1 percentage points), and the impact of most demographic variables is consistent with existing studies. The ordinal quantile models demonstrate significant variation in covariate effects across the distribution of EV purchase intent, offering insights beyond the ordinal probit model. This article is the first to use quantile modeling to reveal how covariate effects differ significantly throughout the spectrum of EV purchase intent.
keywords:
Decarbonization, electric vehicle, ordinal probit, Pew Research, quantile regression, technology adoption.
journal: arXiv.org
1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024). Electric vehicles (EVs) are widely regarded as a key solution for reducing emissions and meeting climate targets, given their zero tailpipe emissions. To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives. In the US, key measures include nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program..
Despite these efforts, EV sales in the US remain relatively low, accounting for only about 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024). At the same time, the future trajectory of EV adoption in the US is increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024). Moreover, successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration has promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its climate agenda, while the Trump administration has signaled a deregulatory shift on EV policies with plans to weaken both incentives as well as regulations. [^2] President Trump’s Executive Order on January 20, 2025, outlined plans to halt \$7.5 billion in charging infrastructure funding, weaken vehicle emissions standards that drive EV production, eliminate EV tax credits, and revoke California’s EV mandate waiver, which would limit state-level regulations. See https://www.whitehouse.gov/presidential-actions/2025/01/unleashing-american-energy/ These shifting political priorities may reflect broader tensions between environmental regulations, industry interests, and fiscal considerations in EV adoption. While EVs help reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024). The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and restrict consumer choice.
Ultimately, the future of the EV market depends on consumer interest, which is likely to be shaped by multiple factors, including cost considerations, convenience, perception of environmental benefits, and awareness of these trade-offs. Although EVs often have higher upfront costs due to expensive battery technology, they can offer long-term savings through lower operational and maintenance expenses, as electric motors require less maintenance. However, the overall cost of owning an EV can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024). Beyond cost considerations, access to reliable charging infrastructure remains an important consideration. Limited charging availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel. Environmental concerns also shape consumer attitudes, but perceptions may vary, given the regional heterogeneity in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
Recognizing the importance of public interest in shaping the future of this market, this study examines how public opinion on purchasing EVs is associated with factors such as exposure to EV information, perceptions about the environmental impact of EVs (as compared to gas-powered vehicles) and government actions to mitigate climate change, confidence in EV related infrastructure development, prior ownership, and demographic characteristics such as age, gender, education, political affiliation, and geographical location of the respondents.
Our paper makes three key contributions. First, we explore the role of several novel factors that have been largely overlooked in the existing EV adoption literature. Previous studies have examined how a combination of factors, including range anxiety, performance concerns, charging infrastructure, environmental attitudes, government policies, peer effects, and demographic characteristics, influences EV purchase intent (Carley et al., 2013; Tiwari et al., 2020; Zhao et al., 2022; Mamkhezri, 2025). Other studies examine the impact of specific factors, such as providing information about total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), reputation (Buhmann and Criado, 2023), and consumer incentives (Xue et al., 2023; Stekelberg and Vance, 2024). Related research examines factors driving actual EV sales rather than purchase intent, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023). We contribute to this literature by examining how EV purchase intent is influenced by factors such as confidence in future charging infrastructure development, perceptions of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits.
Our second contribution is related to usage of large and nationally representative datasets. Existing research studying the trends and determinants of public opinion and purchase intent for EVs has predominantly relied on small-scale surveys. For example, Tiwari et al. (2020) analyzed 1,800 responses from UK residents, Lashari et al. (2021) analyzed 1,500 responses from South Korea, Hu et al. (2023) analyzed 807 responses from China, and Mamkhezri (2025) analyzed 1,500 responses from the US. [^3] While some recent studies (e.g., Pani et al. (2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of perception. Singh et al. (2020) provide a meta-analysis of 211 peer-reviewed research articles published between 2009 and 2019, identifying key factors influencing public opinion of EVs. Notably, most such studies are either based on a single survey for a particular year, not nationally representative, or constrained by sample size; which diminishes their generalizability. This paper overcomes such data related limitations by drawing on three nationally representative opinion polls conducted by the Pew Research Center in the US with a significantly larger sample size than existing previous studies. Specifically, we use data from 2021, 2022, and 2023, with 11,052, 7,173, and 7,201 observations, respectively.
Our third contribution lies in the application of advanced modeling and estimation techniques to study EV purchase intent. Existing research has used various methods to analyze EV purchase intent, such as linear regression to assess the effects of personal beliefs and preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation in shaping EV preferences (Buhmann and Criado, 2023); mixed logit models to explore attitudes towards various EV policy incentives, employment changes and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022). These papers rely on Classical (Frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters. This can lead to biased estimates of covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly translate to changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
In contrast, this paper adopts a Bayesian modeling framework that provides a posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables us to make direct probability statements about the parameters--rather than relying on the notion of “confidence” as in the Classical framework. Our empirical analysis consists of two stages. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how different factors are associated with consumer interest in EV adoption. We present the covariate effects on the average probabilities of outcomes, an aspect which is rarely explored in the EV literature. Next, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how different factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined. Notably, this study is the first to apply ordinal quantile analysis to EV purchase intent. Since no classical estimation methods exist for ordinal quantile models, Bayesian estimation is essential for this analysis. The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
The remainder of the article is structured as follows. Section 2 introduces the ordinal probit and ordinal quantile models employed in our analysis, along with their MCMC algorithms for estimation. Section 3 outlines the data and provides a preliminary investigation. In Sections 4, we present the results from our models and discuss their implications. Finally, Section 5 offers concluding remarks.
2 The Modeling Framework
In this section, we briefly describe the ordinal probit model followed by the ordinal quantile model. For both the models, we outline the Bayesian estimation algorithm using Markov chain Monte Carlo (MCMC) methods and explain the computation of covariate effects.
2.1 The Ordinal Probit Model
Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al. (2023). In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked. However, these scores only carry ordinal meaning and lack cardinal interpretation. Thus, while the categories are ranked, the differences between them cannot be compared directly. For example, in our current study, responses to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
$z_{i}=x^{\prime}_{i}\beta+\varepsilon_{i},\hskip 54.2025pt\forall\;i=1,\cdots,n,$ (1)
where $x_{i}$ is a $k\times 1$ covariate vector including a column of ones, $\beta$ is a $k\times 1$ parameter vector, and $n$ is the number of observations. We assume the error $\varepsilon$ is independent and identically distributed (iid) as a standard normal distribution i.e., $\varepsilon_{i}\sim N(0,1)$ for $i=1,2,\ldots,n$; which gives rise to the ordinal probit model.
In our study, the unobserved variable $z_{i}$ represents intent to purchase EV and is associated with the observed discrete response $y_{i}$, that denotes response to the question about purchasing an EV, as follows:
[…]
where $-\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is visually represented in Figure 1. As seen from the figure, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so on. We set $\gamma_{1}=0$ and $var{(\varepsilon)}=1$ to anchor the location and scale of the distribution, respectively. These restrictions are required for parameter identification and further explained in (Jeliazkov et al., 2008) and Jeliazkov and Rahman (2012).
Figure 1: A pictorial representation of the distribution of the latent variable $z$. The four probabilities P$(y_{i}=1)$, P$(y_{i}=2)$, P$(y_{i}=3)$ and P$(y_{i}=4)$ correspond to the responses: ‘not at all likely’, ‘not too likely’, ‘somewhat likely’, and ‘very likely’, regarding the intention to purchase an electric vehicle ($z_{i}$) for individual $i$ with mean $x^{\prime}_{i}\beta$.
Given the data $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
$\begin{split}f(y|\beta,\gamma)&=\prod_{i=1}^{n}\prod_{j=1}^{J}\Pr(y_{i}=j|\beta,\gamma)^{I(y_{i}=j)}=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}\Phi(\gamma_{j}-x^{\prime}_{i}\beta)-\Phi(\gamma_{j-1}-x^{\prime}_{i}\beta)\bigg{]}^{I(y_{i}=j)},\end{split}$ (3)
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, we apply the transformation of Albert and Chib (2001): $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $12\leq j\leq J-$ 1, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
To complete the Bayesian framework, we utilize the Bayes theorem to combine the augmented likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$. The resulting augmented posterior distribution can be written as,
$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, the correspondence between $\gamma$ and $\delta$ is determined by the one-to-one mapping, and the cutpoint index $j$ is determined by the observed value of $y_{i}$. Additionally, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$ and the priors $\pi(\beta)$ and $\pi(\delta)$ are pdf’s of normal distribution. Based on the augmented joint posterior (4), the conditional posteriors can be derived and the model can be estimated using MCMC sampling, as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effect of a variable. Let’s assume that the $l$-th covariate, ${x_{i,l}}$, is set to two distinct values, $a$ and $b$, denoted as ${x_{i,l}^{a}}$ and ${x_{i,l}^{b}}$, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{p,l},\beta_{p,-l})$, where $-l$ denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). Our goal is to examine the distribution of the difference ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$ for $1\leq j\leq J$, by marginalizing over ${x_{i,-l}}$ and $(\beta_{p},\delta_{p})$. The marginalization enables us to incorporate the uncertainty associated with data and model parameters, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
Algorithm 1: MCMC algorithm for estimating ordinal probit model.
  1. 1.
    Sample $\delta,z|y,\beta$ in one block as follows:
    1. (a)
      Sample $\delta|y,\beta$ marginally of $z$ by generating $\delta^{\prime}$ from a random-walk chain $\delta^{\prime}=\delta+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter, $\hat{D}=\left.-\left[\frac{\partial^{2}\ln f(y|\beta,\delta)}{\partial\delta\partial\delta^{T}}\right]^{-1}\right|_{\delta=\hat{\delta}}$, and $\hat{\delta}=\arg\max_{\delta}\ln f(y|\beta,\delta)$. Accept the proposed value $\delta^{\prime}$ with probability
      $\alpha_{MH}(\delta,\delta^{\prime})=\min\left\{1,\frac{f(y|\beta,\delta^{\prime})\pi(\delta^{\prime})}{f(y|\beta,\delta)\pi(\delta)}\right\},$
      otherwise repeat the current value $\delta$.
    2. (b)
      Sample $z_{i}|y,\beta,\delta\sim TN_{\left(\gamma_{j-1},\gamma_{j}\right)}\left(x_{i}^{\prime}\beta,1\right)$ for $i=1,\ldots,n$, where $TN_{(\gamma_{j-1},\gamma_{j})}$ denotes a truncated normal distribution constrained between $\gamma_{j-1}$ and $\gamma_{j}$; and $\gamma$ is derived from the one-to-one mapping relating $\gamma$ and $\delta$.
  2. 2.
    Sample $\beta|z\sim N(\hat{\beta},\hat{B})$ where $\hat{B}^{-1}=({B}_{0}^{-1}+X^{\prime}X)$, and $\hat{\beta}=\hat{B}(B_{0}^{-1}\beta_{0}+X^{\prime}z)$.
To obtain draws from the distribution ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$, we use the method of composition (Chib and Jeliazkov, 2006). This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference ${\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)}$, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
$ACE_{y=j}\approx\frac{1}{M}\frac{1}{n}\sum_{m=1}^{M}\sum_{i=1}^{n}\Big{[}\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta^{(m)},\delta^{(m)})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta^{(m)},\delta^{(m)})\Big{]},$ (5)
where $(\beta^{(m)},\delta^{(m)})$ denotes the MCMC draws, and $M$ is the number of post burn-in MCMC draws.
2.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression. These insights provide suggestive evidence for more targeted interventions; for example, they show that improving infrastructure visibility may be more effective when directed at consumers already considering EVs, rather than those with little or no interest.
The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
$z_{i}$ $=x^{\prime}_{i}\beta_{p}+\varepsilon_{i},$ $\qquad\forall\;i=1,\cdots,n,$ (6)
[…]
where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005). The AL distribution is essential to create a working likelihood function, and yields the following expression,
$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg{]}^{I(y_{i}=j)},\end{split}$ (7)
where $F_{AL}(\cdot)$ denotes the cdf of an AL distribution and $I(y_{i}=j)$ is an indicator function.
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $12\leq j\leq J-$ 1, where note that $\delta_{p}$ is quantile dependent. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
By Bayes theorem, we combine the augmented likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at the augmented joint posterior distribution,
$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big{\{}\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big{\}}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg{\{}\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg{\}}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$. In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1. In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
The conditional posterior distributions are derived from the augmented joint posterior distribution (i.e. equation 8), and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
  1. 1.
    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
    $\tilde{B}^{-1}_{p}=\bigg{(}\sum_{i=1}^{n}\frac{x_{i}x^{\prime}_{i}}{\tau^{2}w_{i}}+B_{p0}^{-1}\bigg{)},\quad\mbox{and}\quad\tilde{\beta}_{p}=\tilde{B}_{p}\bigg{(}\sum_{i=1}^{n}\frac{x_{i}(z_{i}-\theta w_{i})}{\tau^{2}w_{i}}+B_{p0}^{-1}\beta_{p0}\bigg{)}.$
  2. 2.
    Sample $w_{i}|\beta_{p},z_{i}$ $\sim$ $GIG\,(0.5,\tilde{\lambda}_{i},\tilde{\eta})$, for $i=1,\cdots,n$, where,
    $\tilde{\lambda}_{i}=\Big{(}\frac{z_{i}-x^{\prime}_{i}\beta_{p}}{\tau}\Big{)}^{2},\quad\mbox{and}\quad\tilde{\eta}=\Big{(}\frac{\theta^{2}}{\tau^{2}}+2\Big{)}.$
  3. 3.
    Sample $\delta_{p}|y,\beta_{p}$ marginally of $w$ (latent weight) and $z$ (latent data), by generating $\delta_{p}^{\prime}$ from a random-walk chain $\delta^{\prime}_{p}=\delta_{p}+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter and $\hat{D}$ denotes negative inverse Hessian, obtained by maximizing the log-likelihood with respect to $\delta_{p}$. Given the current value of $\delta_{p}$, accept the proposed draw $\delta^{\prime}_{p}$ with probability,
    $\alpha_{MH}(\delta_{p},\delta^{\prime}_{p})=\min\bigg{\{}1,\frac{\;f(y|\beta_{p},\delta^{\prime}_{p})\;\pi(\beta_{p},\delta^{\prime}_{p})}{f(y|\beta_{p},\delta_{p})\;\pi(\beta_{p},\delta_{p})}\bigg{\}};$
    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
  4. 4.
    Sample $z_{i}|y,\beta_{p},\gamma_{p},w$ $\sim$ $TN_{(\gamma_{p,j-1},\gamma_{p,j})}(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\cdots,n$, where $TN$ denotes a truncated normal distribution and $\gamma_{p}$ is obtained via $\delta_{p}$ using the one-to-one mapping.
After estimating the ordinal quantile models, we compute the covariate effects using the method of composition. Existing studies on quantile regression that have applied this method to estimate the covariate effects include (Rahman and Vossmeyer, 2019) and Bresson et al. (2021). Assuming a setup similar to Section 2.1, our goal is to obtain draws from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. We randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
3 Data
The study draws on survey data from three opinion polls carried out by the Pew Research Center over a three-year span. The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second survey took place during May 2-8, 2022, on a sample of 10,282 adults, and the third survey was administered during May 30-June 4, 2023 on a sample of 10,329 adults. After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis. Table 1 shows a descriptive summary of the independent and dependent variables. Specifically, it presents the mean and standard deviation of the continuous variables, and count and percentage for the categorical variables.
The above survey data offer at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, our survey data is nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage). Second, an appealing feature of our data sets is the availability of ordinal responses to likelihood of purchasing EV. This enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023). Third, we note that the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
Table 1: Descriptive summary of the variables for three Pew Research Center surveys: April 2021; May, 2022; and May-Jun, 2023. Mean (mean) and standard deviations (std) are reported for variables treated as continuous, while count (count) and percentages (per) are reported for categorical variables.
April 2021 May 2022 May–June 2023
variable mean std mean std mean std
Income/10000 (USD) 7.00 3.33 8.00 3.23 8.00 3.24
count per count per count per
EV Info $\circ$ Read or heard a lot 4170 37.73 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ Little or nothing at all 6882 62.27 ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ About the same or worse 3108 28.12 ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident ..absent... . ..absent... . ..absent... . ..absent... . 1263 17.54
$\circ$ Somewhat confident, not too confident, or not at all confident ..absent... . ..absent... . ..absent... . ..absent... . 5938 82.46
EV Owner Owns electric or hybrid vehicle 1039 9.40 769 10.72 812 11.28
GACC $\circ$ Federal government is doing ‘too much’ to reduce the effects of climate change 1962 17.75 1684 23.48 1736 24.11
$\circ$ About the right amount 2203 19.93 1636 22.81 1582 21.97
$\circ$ Too little 6887 62.31 3853 53.72 3883 53.92
Age $<50$ Indicator for age below 50 5124 46.36 2687 37.46 2596 36.05
Female Indicator for female gender 5848 52.91 3819 53.24 3815 52.98
$\circ$ Graduate and above 6447 58.33 3846 53.62 3730 51.80
Education $\circ$ Some college 3194 28.90 2201 30.68 2276 31.61
$\circ$ HS and below 1411 12.77 1126 15.70 1195 16.59
Married Married or living with a partner ..absent... . ..absent... . 5692 79.35 5640 78.32
Metropolitan Lives in a metropolitan area 9817 88.83 6260 87.87 6286 87.29
$\circ$ Northeast 1784 16.14 1100 15.34 1084 15.05
Region $\circ$ Midwest 2444 22.11 1613 22.49 1586 22.02
$\circ$ South 4245 38.41 2841 39.61 2987 41.48
$\circ$ West 2579 23.34 1619 22.57 1544 21.44
US Born Indicator for born in the US 9664 87.44 6150 85.74 6188 85.93
$\circ$ White 7801 70.58 5335 74.38 5083 70.59
Race $\circ$ Black 820 7.42 446 6.22 726 10.08
$\circ$ Other races 2431 22.00 1392 19.41 1392 19.33
$\circ$ Democrat 4201 38.01 2258 31.48 2268 31.50
Party $\circ$ Republican 2667 24.13 2453 34.20 2341 32.51
$\circ$ Independent & others 4184 37.86 2462 34.32 2592 36.00
Very likely 2411 21.82 1526 21.27 1397 19.40
Purchase EV Somewhat likely 2819 25.51 1936 26.99 1815 25.20
Not too likely 3381 30.59 1787 24.91 1736 24.11
Not at all likely 2441 22.09 1924 26.82 2253 31.29
We begin by discussing the independent variables that have a direct influence on purchasing an EV–the question of primary interest in this article. Two such variables, ‘EV Info’ and ‘Env Better’, are specific to the April 2021 survey. The variable ‘EV Info’ measures how much a respondent has read or heard about EVs. It is coded as 1 if the respondent has read or heard a lot about EVs, and 0 if they have read or heard little or nothing. As shown in Table 1, about 38% of respondents have read or heard a lot about EVs. The variable ‘Env Better’ reflects perception about the environmental impact of EVs as compared to gas-powered vehicles. This indicator variable takes the value 1 if the respondent believes that EVs are better for the environment than gas-powered vehicles, and 0 otherwise (worse for the environment or about the same). According to Table 1, approximately 72% of respondents view EVs as being environmentally superior to gas-powered vehicles. Another noteworthy variable is ‘EV Infra’, which is only available in the May-June 2023 survey. This indicator variable is coded as 1 if the respondent is extremely or very confident that the US will build the necessary charging stations and infrastructure to support large number of EVs on the roads, and 0 otherwise (somewhat confident, not too confident, or not at all confident). As seen in Table 1, about 82% of respondents express a negative perception towards the government for creating the infrastructure needed to support the adoption of EVs.
Two key variables of interest to our study, which are available in all the surveys, are ‘EV Owner’ and ‘GACC’. The ‘EV Owner’ variable indicates whether the respondent currently owns an electric or hybrid vehicle and takes the value 1 if true and 0 otherwise. As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets. The ‘GACC’ variable reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC). This variable is divided into three categories: ‘too much’, ‘about the right amount’, and ‘too little’ (with ‘too little’ serving as the reference category). Table 1 indicates that in each sample, more than half of the respondents feel that government efforts to combat global climate change are insufficient.
We also incorporate a broad set of demographic variables as covariates in our regression models. These variables (except marital status) are common across all surveys and include: income, age, gender, education, marital status (not available in the April 2021 survey), metropolitan area, region of residence, birth country, racial background, and party affiliation. We now examine the construction of these variables in more detail.
Pew surveys record income as belonging to one of nine categories: $0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k$, and $>100k$, where $k$ denotes a thousand dollars. To simplify, we use the mid-point of each income range and assign an imputed value of \$20,000 for the first category and \$110,000 for the last category. Age is recorded into four categories (18-29, 30-49, 50-64, and $\geq 65$ years), but for convenience we recode it to a binary indicator taking the value 1 if the respondent is below 50, and 0 otherwise. As shown in Table 1, 46.36% of respondents in the first dataset belong to the younger age group, with this percentage being approximately 10% lower in the following two datasets. Gender is represented by a binary indicator ‘Female’, where 1 indicates female and 0 indicates male. We see that more than 50% respondents across all three datasets are female. Education is classified into three categories: ‘Graduate & Above’, ‘Some College’, and ‘High School (HS) and Below’; the latter is used as the reference category. As shown in Table 1, the majority of respondents are graduates or have higher education, followed by those with some college education. Therefore, respondents across the three samples are generally well-educated.
Information on marital status is available only in the May 2022 and May--June 2023 surveys. We create an indicator variable ‘Married’ that takes value 1 if the respondent is married or living with a partner, and 0 if they are divorced, separated, widowed, or never been married. As per Table 1, more than three-quarters of respondents are married or living with a partner. Two variables reflect a respondent’s residential background. The ‘Metropolitan’ variable is an indicator that equals 1 if the respondent lives in a metropolitan area, and 0 otherwise. Across the three datasets, more than four-fifths of the respondents are located in metropolitan areas. Additionally, we include regional indicators for ‘Northeast’, ‘Midwest’, ‘South’, and ‘West’; with Northeast serving as the reference category in our regression models. As shown in Table 1, the largest proportion of respondents (about 40%) reside in the South, followed by approximately 22% in both the Midwest and West. The smallest fraction, about 15%, lives in the Northeast.
The variable ‘US Born’ indicates whether the respondent was born in the US or is an immigrant. According to Table 1, a predominant fraction (about 86%) of the sample across three datasets were born in the US. Our demographic variable list also includes race indicators for ‘White’ (reference category), ‘Black’, and ‘Other Races’. As expected, the largest proportion of respondents (over 70%) are White, followed by about 20% identifying as belonging to other races. The smallest group, about 8%, are Black. Finally, party affiliation is captured with indicator variables for ‘Democrat’, ‘Republican’ (reference category), and ‘Independent & Others’. Respondents are roughly equally distributed across these three political categories.
As noted earlier, our objective is to analyze the factors that influence individuals’ decision to purchase an EV. However, a continuous measure of support or opposition is not directly observable. The data is limited to the respondent’s answer to the question: “The next time you purchase a vehicle, how likely are you to seriously consider purchasing an electric vehicle?” Response to this question serves as the dependent variable in our models and has four categories: ‘very likely’, ‘somewhat likely’, ‘not too likely’, and ‘not at all likely’. The surveys also include a fifth category :‘I do not expect to purchase a vehicle’; we exclude responses in this category from our analysis to simplify the framework and adhere to ordinal modeling. Table 1 presents the count and percentage of each response category in the three datasets. We see from Table 1, 47.33% of the respondents in the April 2021 survey and 48.26% of the respondents in the May 2022 survey are either ‘very likely’ or ‘somewhat likely’ to purchasing an EV. In comparison, the corresponding percentage in the May-June 2023 survey is 44.60%. Recall, the first two surveys were conducted during the Covid-19 period, while the third survey took place in the post-Covid period. Based on these percentages, we do not observe a significant decline in preference for purchasing EVs in the post-pandemic phase. Overall, our data summary clearly shows that public attitude towards purchasing an EV is generally positive.
Figure 2: Stacked bars displaying the percentage of observations corresponding to the four opinion categories for a selected list of covariates. In each panel, the first stacked bar shows the percentage of observations for ‘not at all likely’ and the cumulative percentage for ‘not at all likely’ and ‘not too likely’. The second stacked bar shows the percentage of observations for ‘somewhat likely’ and the cumulative percentage for ‘somewhat likely’ and ‘very likely’.
Further insights into the association between the dependent variable and selected regressors from the three survey datasets can be gained from Figure 2. In each panel, the bars represent two ‘opinion’ categories: ‘not likely’ and ‘likely’. The first stacked bar displays the percentage of responses for ‘not at all likely’ (shown in dark shade) and the cumulative percentage for both ‘not at all likely’ and ‘not too likely’ (sum of dark and light shades). The second stacked bar illustrates the percentage of responses for ‘somewhat likely’ (in dark shade) and the cumulative percentage for ‘somewhat likely’ and ‘very likely’ (sum of dark and light shades).
The first two panels of Figure 2 highlight the influence of information on consumers’ decisions to purchase EVs. In the first panel, among respondents who have ‘read or heard little or nothing at all’ about EVs, only 46.99% are likely to consider purchasing one. In contrast, the second panel reveals a much higher likelihood of purchase—62.07%—among those who have ‘read or heard a lot’ about EVs. A more striking difference emerges when examining perceptions of the environmental impact of vehicles. As shown in the third panel, just 19.18% of respondents who view EVs as ‘about the same or worse’ than gas-powered vehicles are likely to purchase one. However, as shown in the fourth panel, the percentage jumps to 65.79% for those who believe EVs have a lower environmental impact. A similar pattern is evident, from panels five and six, when considering respondents’ views on the development of charging stations and infrastructure to support large number of EVs on the road. Among those who lack confidence that the government will build the necessary facilities, only 36.49% are likely to buy an EV. In stark contrast, a much higher 82.74% of respondents who believe the government will build the required facilities for EVs are likely to make a purchase. These relationships highlight how knowledge and environmental impact of EVs, and confidence in supporting infrastructure play pivotal roles in shaping consumer purchase decisions.
The remainder of panels in Figure 2 (i.e., panels seven to eighteen) exhibit how factors such as ownership of an EV or hybrid vehicle, age, and gender affect the decision to purchase an EV. The stacked bars from panels seven to twelve clearly show that owning an EV or hybrid vehicle (as opposed to not) increases the likelihood of purchasing an EV across all survey datasets. For instance, 84.12% of EV or hybrid owners are likely to purchase an EV as compared to only 49.42% (a difference of 34.7%) in the April 2021 survey data. For the same comparison, a difference of 36.98% ($81.28\%-44.3\%$) and 36.47% ($76.97\%-40.5\%$) is noted in the May 2022 and May-June 2023 survey data. Age also plays a role, with a majority of respondents under 50 being likely to purchase an EV across all three datasets. With respect to gender, the likelihood of purchasing an EV among females is 53.27% in the April 2021 dataset, decreases to 49.18% in May 2022, and further drops to 44.69% in the May-June 2023 dataset.
4 Empirical Results and Discussion
In this section, we report the results from the Bayesian estimation of ordinal probit and ordinal quantile models. We also report the marginal likelihood which may be used for model comparison. Morevoer, for each model, we compute the covariate effects for all or some of the significant variables and discuss our findings in the context of existing literature.
4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey data. Specifically, Table 2 presents the posterior means and standard deviations of the parameters along with the logarithm of marginal likelihood (Chib and Jeliazkov, 2001). The estimates are obtained from 10,000 MCMC iterations (after a burn-in of 2,500 iterations) and utilizes the following prior distributions: $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\ast I)$, where $I$ denotes an identity covariance matrix. The priors are relatively diffuse to allow the data speak for itself. In ordinal models, the sign of coefficients indicates the direction of covariate effects for the first and last response categories. For example, in Table 2, the coefficient for “EV Info” is positive (0.43), which implies that it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”) to purchase an EV. However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the covariate effects on intermediate outcomes (i.e., “Not too likely” and “Somewhat likely”). Therefore, Table 3 provides the covariate effects for all variables which are statistically different from zero. Note, our measure of covariate effects account for uncertainty in parameters and observations to avoid biases (Jeliazkov and Vossmeyer, 2018).
The combined results from the 2021, 2022, and 2023 surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects corresponding to the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs as compared to those with limited exposure to EV-related information. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings align with Zhao et al. (2022) and Scherrer (2023), who demonstrate that greater exposure to information about EVs—whether through peer effects or media—positively influences attitudes toward EVs, and with Krause et al. (2013), who underscore the crucial role of accurate information in shaping consumer perceptions.
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, we also observe, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding aligns with Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs  [^4] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood, from April 2021; May 2022; and May–June 2023 survey data.
April 2021 May 2022 May–June 2023
Mean Std Mean Std Mean Std
Intercept $-0.28$ $0.07$ $0.36$ $0.09$ $-0.05$ $0.09$
EV Info $0.43$ $0.02$.absent... . ..absent... . ..absent... . ..absent... .
Env Better $0.79$ $0.03$.absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . $0.82$ $0.04$
EV Owner $0.79$ $0.04$ $0.86$ $0.05$ $0.82$ $0.04$
GACC (TM) $-0.79$ $0.04$ $-1.13$ $0.04$ $-1.07$ $0.04$
GACC (RA) $-0.52$ $0.03$ $-0.63$ $0.03$ $-0.49$ $0.04$
Age $<50$ $0.16$ $0.02$ $0.17$ $0.03$ $0.15$ $0.03$
Female $-0.02$ $0.02$ $-0.11$ $0.03$ $-0.12$ $0.03$
Grad & Above $0.38$ $0.04$ $0.45$ $0.04$ $0.49$ $0.04$
Some College $0.18$ $0.04$ $0.23$ $0.04$ $0.21$ $0.04$
Married ..absent... . ..absent... . $0.06$ $0.03$ $0.02$ $0.04$
Metropolitan $0.24$ $0.03$ $0.25$ $0.04$ $0.20$ $0.04$
Midwest $-0.08$ $0.03$ $-0.05$ $0.04$ $-0.07$ $0.04$
South $-0.05$ $0.03$ $-0.02$ $0.04$ $-0.07$ $0.04$
West $0.19$ $0.03$ $0.19$ $0.04$ $0.18$ $0.04$
US Born $-0.21$ $0.04$ $-0.33$ $0.04$ $-0.22$ $0.04$
Black $-0.28$ $0.04$ $-0.14$ $0.06$ $-0.18$ $0.05$
Other Races $0.05$ $0.03$ $0.01$ $0.04$ $-0.03$ $0.04$
Democrat $0.59$ $0.03$ $0.65$ $0.04$ $0.75$ $0.04$
Ind & Others $0.40$ $0.03$ $0.36$ $0.03$ $0.43$ $0.03$
Income/10000 (USD) $0.01$ $0.01$ $0.02$ $0.01$ $0.03$ $0.01$
$\delta_{1}$ $0.01$ $0.02$ $-0.10$ $0.02$ $-0.11$ $0.02$
$\delta_{2}$ $0.12$ $0.01$ $-0.00$ $0.02$ $0.04$ $0.02$
$\ln$ ML $-12415$ $-8301$ $-7914$
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents who are very or extremely confident that the US will build the necessary charging infrastructure are 19.1 percentage points more likely to express strong interest in purchasing EVs and 18.8 percentage points less likely to report no interest compared to those who report lower confidence. This finding aligns with prior research, such as Carley et al. (2013) and Li et al. (2017). As shown in Table 1, 82% of respondents lack confidence that US will develop the infrastructure needed to support EV adoption. Taken together, these results suggest that skepticism about charging infrastructure is likely to be a significant barrier to broader adoption, highlighting the importance of both improving infrastructure and clearly communicating its availability, reliability, and expansion plans to the public.
Table 3: Covariate effects (CE) from ordinal probit models. CE for variables not statistically different from zero are marked with $\times$.
April 2021 May 2022 May–June 2023
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0851$ $-0.0388$ $0.0217$ $0.1022$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $-0.1955$ $-0.0565$ $0.0970$ $0.1551$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $-0.1880$ $-0.0553$ $0.0526$ $0.1908$
EV Owner $-0.1321$ $-0.0849$ $0.0051$ $0.2119$ $-0.1749$ $-0.0794$ $0.0231$ $0.2312$ $-0.1814$ $-0.0513$ $0.0421$ $0.1905$
GACC (TM) $0.1971$ $0.0453$ $-0.0945$ $-0.1478$ $0.3341$ $0.0133$ $-0.1507$ $-0.1968$ $0.3170$ $-0.0153$ $-0.1395$ $-0.1622$
GACC (RA) $0.1180$ $0.0391$ $-0.0490$ $-0.1082$ $0.1649$ $0.0267$ $-0.0639$ $-0.1278$ $0.1282$ $0.0079$ $-0.0486$ $-0.0875$
Age $<50$ $-0.0326$ $-0.0130$ $0.0099$ $0.0356$ $-0.0410$ $-0.0090$ $0.0129$ $0.0371$ $-0.0371$ $-0.0036$ $0.0126$ $0.0282$
Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ $-0.0082$ $-0.0240$ $0.0301$ $0.0027$ $-0.0100$ $-0.0228$
Grad & Above $-0.0815$ $-0.0322$ $0.0276$ $0.0860$ $-0.1139$ $-0.0240$ $0.0378$ $0.1001$ $-0.1264$ $-0.0130$ $0.0445$ $0.0949$
Some College $-0.0373$ $-0.0152$ $0.0099$ $0.0426$ $-0.0541$ $-0.0118$ $0.0156$ $0.0503$ $-0.0517$ $-0.0052$ $-0.0166$ $0.0403$
Married ..absent... . ..absent... . ..absent... . ..absent... . $-0.0154$ $-0.0029$ $0.0050$ $0.0134$$\times$ $\times$ $\times$ $\times$
Metropolitan $-0.0534$ $-0.0179$ $0.0191$ $0.0522$ $-0.0662$ $-0.0099$ $0.0232$ $0.0528$ $-0.0512$ $-0.0027$ $0.0183$ $0.0356$
Midwest $0.0162$ $0.0061$ $-0.0051$ $-0.0172$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
South $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West $-0.0378$ $-0.0164$ $0.0101$ $0.0441$ $-0.0457$ $-0.0109$ $0.0135$ $0.0431$ $-0.0454$ $-0.0053$ $0.0148$ $0.0360$
US Born $0.0416$ $0.0186$ $-0.0106$ $-0.0497$ $0.0777$ $0.0217$ $-0.0214$ $-0.0781$ $0.0551$ $0.0071$ $-0.0178$ $-0.0444$
Black $0.0614$ $0.0200$ $-0.0220$ $-0.0594$ $0.0349$ $0.0058$ $-0.0116$ $-0.0291$ $0.0460$ $0.0026$ $-0.0160$ $-0.0326$
Other Races $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat $-0.1211$ $-0.0578$ $0.0403$ $0.1386$ $-0.1569$ $-0.0497$ $0.0513$ $0.1553$ $-0.1882$ $-0.0353$ $0.0677$ $0.1558$
Ind & Others $-0.0808$ $-0.0310$ $0.0219$ $0.0898$ $-0.0890$ $-0.0187$ $0.0266$ $0.0811$ $-0.1091$ $-0.0093$ $0.0356$ $0.0827$
$\Delta$I=\$$=\$50,000$$\times$ $\times$ $\times$ $\times$ $-0.0223$ $-0.0051$ $0.0066$ $0.0209$ $-0.0363$ $-0.0045$ $0.0116$ $0.0291$
Across all three surveys, previous EV ownership is consistently linked with a higher interest in purchasing an EV. Respondents who have owned an EV or hybrid vehicle are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest compared to those without prior ownership. Other studies such as Carley et al. (2013) and Buhmann and Criado (2023) have also documented this pattern, suggesting that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in the technology, making them more engaged with the EV market.
Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest. Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though the effects are smaller.
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs compared to older respondents. Our finding aligns with Zhao et al. (2022), Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025). For instance, Zhao et al. (2022) finds that EV buyers in Shanghai are predominantly younger, with most between the ages of 31 and 40.
Results from the 2022 and 2023 surveys also show that female respondents are modestly less likely to express strong interest and more likely to report no interest than males. The covariate effects on strong interest decrease between 2.3–2.4 percentage points; while the covariate effect for no interest increase between 2.7–3.0 percentage points. This finding is consistent several studies, including Carley et al. (2013), Krause et al. (2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men in their samples exhibit a stronger preference for EVs.
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest compared to those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. The positive relationship between education and EV interest is similarly documented in Carley et al. (2013), Krause et al. (2013), Buhmann and Criado (2023), and Mamkhezri (2025).
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report EV preference to be higher among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest compared to those in the Northeast, though those in the Midwest and South show no significant difference from their Northeastern counterparts. This pattern likely reflects California’s leadership in EV adoption, driven in part by policies such as ZEV mandates and the Clean Vehicle Rebate Program, along with a robust charging network. [^5] See https://www.eia.gov/todayinenergy/detail.php?id=61082
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest compared to non-US-born respondents, possibly reflecting differences in cultural attitudes, driving patterns, or perceptions about the cost of vehicle ownership. Such a difference in perception about EVs is also evident in the racial background of respondents. Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents, likely reflecting structural barriers, such as limited access to charging infrastructure (Hsu and Fingerman, 2021). Other non-white respondents show similar preferences to White respondents.
Political affiliation is a consistent and significant predictor of EV purchase intent. Democrats are 13.9–18.8 percentage points more likely to report strong interest in EVs and 12.1–18.8 percentage points less likely to report no interest compared to Republicans. Similarly, Independents and others are also about 8 percentage points more likely to express strong interest in purchasing EVs relative to Republicans. These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible compared to Republicans, as documented by Sintov et al. (2020). They also reflect the starkly contrasting policy approaches of the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), President Trump’s administration rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In contrast, the Biden administration adopted an aggressive pro-EV stance as part of its broader climate agenda, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
Conditional on other factors, an additional \$50,000 in income is either not significantly associated with EV interest or changes the probability of strong interest (no interest) by a meagre 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals. The small covariate effect for income may reflect the growing availability of EVs across various market segments and the presence of subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
Overall, the results show that interest in EVs varies based on environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographic factors. Some patterns likely reflect personal preferences, while others may reflect barriers such as limited awareness or charging access.
4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to the probabilities of outcomes at specific quantiles, conditional on the covariates. In particular, Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models along with the logarithm of marginal likelihood, estimated separately for each survey data. Table 5 reports the corresponding covariate effects. The covariate effects at the 50th percentile align closely with those from the ordinal probit model. However, across all surveys, the covariate effects at the 20th and 80th quantiles are noticeably different compared to those from the ordinal probit model. As such, we find compelling evidence that the impact of covariates on the intent to purchase EV is not uniform, but varies across the distribution of EV purchase intent. This highlights the value of quantile modeling for capturing these variations. Below, we emphasize the key differences across quantiles.
While having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption. In the 2021 survey and at the 80th percentile of EV purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent to purchase compared to those who have read or heard little to nothing. At the 50th and 20th percentiles of EV purchase intent, the difference is 9.9 and 6.1 percentage points, respectively. The smaller covariate effect of prior exposure to EV information at the 20th percentile, suggests that simply raising general awareness may not be enough to persuade consumers with low purchase interest. In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may be effective in influencing purchase decisions.
Similarly, the relationship between confidence that the US will build the necessary charging infrastructure (measured in the 2023 survey) and EV purchase intent varies across the distribution of EV purchase intent. At the 80th percentile, respondents who are confident in this development are 20.0 percentage points more likely to express strong intent to purchase compared to those who are not confident. The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile. Conversely, at the 80th percentile, those confident in infrastructure development (relative to those who are not) are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively. These findings suggest that confidence in infrastructure development strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent. Enhancing the affordability, accessibility, and reliability of charging infrastructure—and effectively communicating these efforts—can build consumer confidence and promote broader EV adoption.
Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood, from April 2021, May 2022, and May–Jun 2023 survey data.
April 2021 May 2022 May–June 2023
20th 50th 80th 20th 50th 80th 20th 50th 80th
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
Intercept $-3.60$ $0.25$ $-0.58$ $0.17$ $2.60$ $0.25$ $-2.15$ $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ $-2.81$ $0.30$ $-0.05$ $0.20$ $2.77$ $0.28$
EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
EV Owner $2.80$ $0.13$ $1.76$ $0.10$ $2.91$ $0.16$ $2.94$ $0.16$ $1.83$ $0.11$ $3.31$ $0.20$ $2.65$ $0.16$ $1.99$ $0.11$ $3.08$ $0.19$
GACC (TM) $-2.81$ $0.13$ $-1.96$ $0.09$ $-2.49$ $0.13$ $-3.79$ $0.16$ $-2.63$ $0.10$ $-3.70$ $0.14$ $-4.00$ $0.17$ $-2.41$ $0.10$ $-3.45$ $0.14$
GACC (RA) $-1.79$ $0.10$ $-1.24$ $0.07$ $-1.55$ $0.11$ $-1.94$ $0.12$ $-1.48$ $0.08$ $-2.03$ $0.12$ $-1.77$ $0.12$ $-1.15$ $0.08$ $-1.48$ $0.12$
Age$<50$ $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
Female $-0.11$ $0.08$ $-0.02$ $0.05$ $-0.03$ $0.07$ $-0.36$ $0.10$ $-0.22$ $0.06$ $-0.33$ $0.09$ $-0.49$ $0.10$ $-0.31$ $0.07$ $-0.27$ $0.09$
Grad & Above $1.42$ $0.14$ $-0.93$ $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
Some College $0.66$ $0.14$ $0.47$ $0.09$ $0.57$ $0.12$ $1.08$ $0.16$ $0.53$ $0.10$ $0.49$ $0.13$ $0.92$ $0.16$ $0.42$ $0.10$ $0.53$ $0.13$
Married ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
Metropolitan $0.77$ $0.12$ $0.64$ $0.08$ $0.71$ $0.12$ $0.91$ $0.16$ $0.59$ $0.10$ $0.79$ $0.13$ $0.79$ $0.16$ $0.46$ $0.10$ $0.61$ $0.13$
Midwest $-0.31$ $0.12$ $-0.19$ $0.08$ $-0.24$ $0.11$ $-0.09$ $0.15$ $-0.09$ $0.10$ $-0.11$ $0.14$ $-0.18$ $0.16$ $-0.15$ $0.10$ $-0.21$ $0.14$
South $-0.21$ $0.10$ $-0.12$ $0.07$ $-0.17$ $0.10$ $0.00$ $0.13$ $-0.02$ $0.09$ $-0.06$ $0.13$ $-0.22$ $0.15$ $-0.22$ $0.09$ $-0.17$ $0.13$
West $0.57$ $0.12$ $0.47$ $0.08$ $0.70$ $0.12$ $0.46$ $0.15$ $0.45$ $0.10$ $0.89$ $0.14$ $0.63$ $0.16$ $0.46$ $0.10$ $0.66$ $0.15$
US Born $-0.63$ $0.12$ $-0.49$ $0.09$ $-0.80$ $0.14$ $-0.89$ $0.15$ $-0.67$ $0.10$ $-1.25$ $0.16$ $-0.91$ $0.16$ $-0.45$ $0.11$ $-0.67$ $0.15$
Black $-1.23$ $0.15$ $-0.70$ $0.10$ $-0.56$ $0.14$ $-0.86$ $0.20$ $-0.25$ $0.14$ $-0.06$ $0.18$ $-0.94$ $0.16$ $-0.42$ $0.11$ $-0.36$ $0.16$
Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ $-0.19$ $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ $-0.30$ $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
Democrat $2.00$ $0.11$ $1.33$ $0.08$ $1.94$ $0.12$ $2.30$ $0.14$ $1.45$ $0.09$ $1.90$ $0.13$ $2.49$ $0.14$ $1.71$ $0.10$ $2.40$ $0.14$
Ind & Others $1.29$ $0.10$ $0.88$ $0.07$ $1.42$ $0.11$ $1.32$ $0.12$ $0.80$ $0.08$ $1.06$ $0.11$ $1.40$ $0.13$ $0.95$ $0.08$ $1.44$ $0.12$
Income/10000 (USD) $0.04$ $0.01$ $0.02$ $0.01$ $0.00$ $0.01$ $0.07$ $0.02$ $0.05$ $0.01$ $0.05$ $0.02$ $0.12$ $0.02$ $0.06$ $0.01$ $0.08$ $0.02$
$\delta_{1}$ $1.16$ $0.02$ $0.91$ $0.02$ $1.45$ $0.02$ $1.09$ $0.02$ $0.76$ $0.02$ $1.25$ $0.02$ $1.14$ $0.02$ $0.74$ $0.02$ $1.21$ $0.02$
$\delta_{2}$ $1.53$ $0.02$ $0.95$ $0.02$ $1.35$ $0.02$ $1.46$ $0.02$ $0.84$ $0.02$ $1.16$ $0.02$ $1.52$ $0.02$ $0.93$ $0.02$ $1.19$ $0.02$
$\ln$ ML $-12871$ $-12529$ $-12901$ $-8563$ $-8357$ $-8624$ $-8237$ $-7981$ $-8194$
Table 5: Covariate effects (CE) from quantile analysis. CE for variables not statistically different from zero are marked with $\times$.
April 2021 Survey
20th Quantile 50th Quantile 80th Quantile
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0722$ $-0.0421$ $0.0527$ $0.0615$ $-0.0792$ $-0.0374$ $0.0172$ $0.0995$ $-0.0722$ $-0.0453$ $-0.0089$ $0.1265$
Env Better $-0.1958$ $-0.0191$ $0.1135$ $0.1014$ $-0.1771$ $-0.0730$ $0.1122$ $0.1379$ $-0.1384$ $-0.1127$ $0.0789$ $0.1722$
EV Owner $-0.1357$ $-0.0896$ $0.0633$ $0.1620$ $-0.1212$ $-0.0757$ $-0.0113$ $0.2083$ $-0.1053$ $-0.0953$ $-0.0265$ $0.2270$
GACC (TM) $0.2192$ $0.0097$ $-0.1217$ $-0.1072$ $0.1974$ $0.0635$ $-0.1203$ $-0.1406$ $0.1283$ $0.0779$ $-0.0575$ $-0.1487$
GACC (RA) $0.1216$ $0.0312$ $-0.0780$ $-0.0747$ $0.1097$ $0.0534$ $-0.0613$ $-0.1018$ $0.0763$ $0.0421$ $-0.0124$ $-0.1059$
Age $<50$ $-0.0237$ $-0.0106$ $0.0156$ $0.0187$ $-0.0303$ $-0.0134$ $0.0095$ $0.0343$ $-0.0265$ $-0.0164$ $0.0003$ $0.0427$
Metropolitan $-0.0483$ $-0.0174$ $0.0318$ $0.0338$ $-0.0565$ $-0.0225$ $0.0237$ $0.0553$ $-0.0345$ $-0.0185$ $0.0030$ $0.0500$
Midwest $0.0185$ $0.0080$ $-0.0123$ $-0.0141$ $0.0161$ $0.0068$ $-0.0053$ $-0.0177$ $0.0114$ $0.0066$ $-0.0003$ $-0.0177$
South $0.0124$ $0.0056$ $-0.0083$ $-0.0098$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West $-0.0327$ $-0.0168$ $0.0220$ $0.0275$ $-0.0382$ $-0.0176$ $0.0092$ $0.0467$ $-0.0311$ $-0.0200$ $-0.0015$ $0.0526$
Black $0.0784$ $0.0256$ $-0.0519$ $-0.0520$ $0.0613$ $0.0244$ $-0.0259$ $-0.0598$ $0.0269$ $0.0141$ $-0.0014$ $-0.0396$
Democrat $-0.1242$ $-0.0523$ $0.0808$ $0.0958$ $-0.1081$ $-0.0631$ $0.0431$ $0.1280$ $-0.0828$ $-0.0676$ $0.0070$ $0.1434$
May 2022 Survey
EV Owner $-0.1782$ $-0.0723$ $0.0880$ $0.1624$ $-0.1573$ $-0.0695$ $0.0121$ $0.2147$ $-0.1465$ $-0.1110$ $-0.0071$ $0.2646$
GACC (TM) $0.3425$ $-0.0568$ $-0.1515$ $-0.1341$ $0.3304$ $0.0226$ $-0.1778$ $-0.1752$ $0.2471$ $0.0963$ $-0.1459$ $-0.1975$
GACC (RA) $0.1541$ $0.0026$ $-0.0790$ $-0.0777$ $0.1527$ $0.0497$ $-0.0852$ $-0.1171$ $0.1269$ $0.0392$ $-0.0308$ $-0.1352$
Age $<50$ $-0.0362$ $-0.0066$ $0.0199$ $0.0229$ $-0.0390$ $-0.0085$ $0.0130$ $0.0344$ $-0.0358$ $-0.0167$ $0.0072$ $0.0453$
Metropolitan $-0.0684$ $-0.0054$ $0.0360$ $0.0378$ $-0.0643$ $-0.0109$ $0.0258$ $0.0493$ $-0.0483$ $-0.0165$ $0.0114$ $0.0535$
Black $0.0642$ $0.0060$ $-0.0342$ $-0.0359$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat $-0.1719$ $-0.0295$ $0.0927$ $0.1087$ $-0.1437$ $-0.0574$ $0.0601$ $0.1410$ $-0.0996$ $-0.0659$ $0.0217$ $0.1438$
May–June 2023 Survey
EV Infra $-0.1863$ $-0.0428$ $0.0956$ $0.1335$ $-0.1827$ $-0.0603$ $0.0616$ $0.1813$ $-0.1452$ $-0.0730$ $0.0187$ $0.1995$
EV Owner $-0.1684$ $-0.0350$ $0.0727$ $0.1307$ $-0.1842$ $-0.0542$ $0.0459$ $0.1925$ $-0.1580$ $-0.0817$ $0.0183$ $0.2215$
GACC (TM) $0.3323$ $-0.0635$ $-0.1405$ $-0.1283$ $0.3103$ $-0.0250$ $-0.1426$ $-0.1427$ $0.2512$ $0.0486$ $-0.1423$ $-0.1575$
GACC (RA) $0.1318$ $-0.0065$ $-0.0592$ $-0.0661$ $0.1238$ $0.0132$ $-0.0570$ $-0.0799$ $0.0975$ $0.0183$ $-0.0329$ $-0.0830$
Age $<50$ $-0.0296$ $-0.0007$ $0.0131$ $0.0173$ $-0.0363$ $-0.0034$ $0.0139$ $0.0258$ $-0.0349$ $-0.0082$ $0.0102$ $0.0329$
Metropolitan $-0.0568$ $0.0013$ $0.0250$ $0.0304$ $-0.0511$ $-0.0026$ $0.0204$ $0.0333$ $-0.0408$ $-0.0067$ $0.0129$ $0.0345$
Black $0.0665$ $-0.0010$ $-0.0296$ $-0.0359$ $0.0462$ $0.0029$ $-0.0182$ $-0.0309$ $0.0239$ $0.0039$ $-0.0069$ $-0.0209$
Democrat $-0.1832$ $-0.0084$ $0.0851$ $0.1066$ $-0.1815$ $-0.0384$ $0.0798$ $0.1401$ $-0.1393$ $-0.0660$ $0.0491$ $0.1562$
Beliefs about EVs’ environmental benefits (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles. Among respondents at the 80th percentile, those who believe EVs are better for the environment are 17.2 percentage points more likely to express strong purchase intent compared to those at the same percentile who do not share this belief. The corresponding differences are 13.8 percentage points at the median and 10.1 percentage points at the 20th percentile. Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 percentage points less likely at the median and 19.6 percentage points less likely at the 20th percentile. This finding suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters while easing skepticism among consumers with lower purchase interest; this indicates that highlighting benefits such as local air quality improvements could broaden support for EV adoption.
The influence of prior EV ownership also varies across quantiles. At the 80th percentile, respondents with prior EV ownership are 22.7 percentage points more likely to express strong purchase intent compared to those without prior ownership. This difference decreases to 20.8 percentage points at the median and 16.2 percentage points at the 20th percentile. A similar pattern is observed in both the 2022 and 2023 surveys. In the 2022 survey, the probability of strong interest increases by 26.5, 21.5, and 16.2 percentage points at the 80th, 50th, and 20th quantiles, respectively. For the same quantiles, the covariate effects for strong interest in the 2023 survey are 22.1, 19.2, and 13.1 percentage points.
When it comes to views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong interest in purchasing an EV, with this negative effect becoming more pronounced at higher percentiles. In the 2021 survey, those at the 80th percentile of EV purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent compared to those at the same percentile who do not. The gap is 14.1 percentage points at the median and 10.7 percentage points at the 20th percentile. Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 percentage points at the median and 21.9 percentage points at the 20th percentile. More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much on climate issues being approximately 33 percentage points more likely to express no interest in EVs at the 20th percentile and 25 percentage points more likely at the 80th percentile. These patterns suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonates across ideological lines. Rather than positioning EVs primarily as a climate solution, emphasizing other benefits—such as long-term cost savings and air quality improvements—may be more effective in engaging a broader range of potential buyers. Moreover, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns, particularly for consumers who are wary of climate-focused policies.
Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments. Legend: ΔPr(Not at all $\Delta\Pr(\textrm{Not at all likely})$:  ; ΔPr(Not too $\Delta\Pr(\textrm{Not too likely})$:   ; ΔPr(Somewhat $\Delta\Pr(\textrm{Somewhat likely})$:   ; ΔPr(Very $\Delta\Pr(\textrm{Very likely})$:  .
Geographic and demographic factors show smaller but consistent relationships with EV purchase intent across all surveys. Younger respondents (under 50), metropolitan residents, and those living in the West are more likely to express higher purchase intent across all considered quantiles, with the strongest covariate effects observed at the 80th percentile.
Finally, the relationship between political party affiliation and EV purchase intent varies considerably across quantiles. Results from the 2021 survey show that at the 80th percentile Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points among respondents at the 20th percentile. Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, while the difference was 12.4 percentage points at the 20th percentile. These patterns remain consistent in more recent surveys. The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, while at lower quantiles, they are far less likely to reject EVs outright. Emphasizing benefits that resonate across political ideologies—such as long-term cost savings and energy independence—may help broaden EV appeal across the political spectrum.
To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents the covariate effects for some selected variables from a series of ordinal quantile regression models, covering the 5th to the 95th percentiles of EV purchase intent in 5-percentile increments. Each panel corresponds to a different covariate; with the solid gray, dashed gray, dashed black, and solid black lines representing the change in probability of reporting “not likely at all,” “not too likely,” “somewhat likely,” and “very likely,” respectively, for a given change in covariate across the quantiles. The figure clearly illustrates the substantial differences in covariate effects across the distribution of EV purchase intent.
For example, the solid black line in these panels indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have previously owned an EV, identify as Democrats, or are confident in the United States’ ability to develop the necessary infrastructure are much more likely to report strong interest in purchasing EVs compared to those at lower quantiles. Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest in buying EVs at lower quantiles and markedly less likely to report strong interest in buying EVs at higher quantiles.
Overall, these results highlight the need for tailored strategies to address the barriers to EV adoption for different population segments.
5 Conclusion
EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts. However, government support has varied between administrations. The Obama administration facilitated EV adoption by introducing federal tax credits for EV purchases and strengthening fuel economy and emissions standards, which incentivized automakers to develop electric models. In contrast, the Trump administration’s first term prioritized deregulation and rolled back these standards. The Biden administration implemented stricter emissions regulations and allocated substantial funding for EV infrastructure through the Bipartisan Infrastructure Law. A second Trump administration has signaled a likely return to deregulation, though specific policies remain unclear.
While policies and regulatory standards continue to evolve with changing administrations, the future of electric vehicles depends largely on the interest of consumers, which plays a crucial role in shaping national policies. Public opinion has a well-documented influence on US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has systematically analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of the covariates. In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
This study addresses these gaps by examining public attitudes toward EVs using three nationally representative surveys conducted by the Pew Research Center from 2021-2023. Given the ordinal nature of the response data, we employ both ordinal probit regression and ordinal quantile regression to model EV purchase intent as a function of four novel variables not considered in existing studies, prior EV ownership, political affiliation, and demographic characteristics such as age, gender, income, and geographic location. We adopt the Bayesian approach to modeling and estimation, as it offers the entire posterior distribution of the parameters, enables direct probability statements, and allows us to account for uncertainty in observations and parameters when evaluating the effects of covariates.
The results from ordinal probit models indicate that respondents who say they have read or heard a lot about EVs, perceive them as environmentally beneficial, have previously owned an EV or hybrid vehicle, and are confident in the US developing charging infrastructure are more likely to express strong interest in purchasing an EV. Other factors positively influencing the intention to purchase an EV include being under 50 years old, having a higher level of education, living in metropolitan areas, residing in the West (compared to the Northeast), identifying with the Democratic Party (rather than the Republican Party), and having a higher income. In contrast, those who believe government climate efforts are excessive or adequate, are female, live in the Midwest, were born in the US, or are Black (as opposed to White) are less likely to express strong interest and more likely to report no interest in EVs. Results from covariate effects reveal that prior EV ownership, confidence in EV infrastructure, belief in EVs’ environmental benefits, attitudes toward government climate policy, and political affiliation have the strongest influence on individuals’ decision to purchase EVs.
In addition, results from ordinal quantile analysis show that the relationship between these factors and EV purchase intent varies across the conditional distribution of the intention to purchase EVs. Respondents who are knowledgeable about EVs, believe in their environmental benefits, are confident in US charging infrastructure, have prior EV ownership, or identify as Democrats are especially likely to express strong purchase intent at higher quantiles of EV purchase intent. Similarly, those who believe EVs are environmentally beneficial, have confidence in infrastructure development, or identify as Democrats are particularly less likely to express no interest at lower quantiles. In contrast, respondents who believe the government is doing too much to address climate change are particularly more likely to express no interest at lower quantiles and particularly less likely to express strong interest at higher quantiles. Overall, the quantile models highlight significant variation in covariate effects across the distribution of EV purchase intent; thereby providing a more nuanced understanding than the ordinal probit model.
The findings demonstrate how targeted strategies may be able to help accelerate EV adoption. For consumers with low purchase intent (i.e., at lower quantiles), general awareness campaigns may have limited impact; instead, emphasizing practical aspects such as long-term cost savings, availability and reliability of charging infrastructure, and financial incentives could be more effective. For more interested consumers (i.e., at higher quantiles), reinforcing beliefs about environmental benefits can transform a likely customer to an actual EV owner. Confidence in charging infrastructure plays a dual role, both strengthening commitment among likely adopters and reducing outright rejection among skeptics, highlighting the potential value of visible investments and clear public communication about infrastructure progress. The results also indicate that political identity and views on climate policy are closely related to EV purchase intent, suggesting that framing EVs in terms of widely appealing features such as financial savings and energy independence—rather than solely as a climate solution—could resonate with a broader range of potential buyers. Moreover, for consumers wary of government climate policy, information provided by automakers, dealerships, and industry groups may carry more credibility than government-led campaigns. Overall, the empirical findings of our study provide useful insights for policymakers, industry stakeholders, and advertisers seeking to bolster adoption.

### v2

@@
-To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States
+Do Determinants of EV Purchase Intent vary across the Spectrum? Evidence from Bayesian Analysis of US Survey Data
 Abstract
-The adoption of electric vehicles (EVs) is considered critical to achieving climate goals, yet it hinges on consumer interest.
-This study explores how public intent to purchase EVs relates to four unexamined factors: exposure to EV information, perceptions of EVs’ environmental benefits, views on government climate policy, and confidence in future EV infrastructure; while controlling for prior EV ownership, political affiliation, and demographic characteristics (e.g., age, gender, education, and geographic location).
-We utilize data from three nationally representative opinion polls conducted by the Pew Research Center between 2021 and 2023, and employ Bayesian techniques to estimate the ordinal probit and ordinal quantile models.
-Results from ordinal probit show that respondents who are well-informed about EVs, perceive them as environmentally beneficial, or are confident in development of charging stations are more likely to express strong interest in buying an EV, with covariate effects--a metric rarely reported in EV research--of 10.2, 15.5, and 19.1 percentage points, respectively.
-In contrast, those skeptical of government climate initiatives are more likely to express no interest, by more than 10 percentage points.
-Prior EV ownership exhibits the highest covariate effect (ranging from 19.0 to 23.1 percentage points), and the impact of most demographic variables is consistent with existing studies.
-The ordinal quantile models demonstrate significant variation in covariate effects across the distribution of EV purchase intent, offering insights beyond the ordinal probit model.
-This article is the first to use quantile modeling to reveal how covariate effects differ significantly throughout the spectrum of EV purchase intent.
+While electric vehicle (EV) adoption has been widely studied, most research focuses on the average effects of predictors on purchase intent, overlooking variation across the distribution of EV purchase intent.
+This paper makes a threefold contribution by analyzing four unique explanatory variables, leveraging large-scale US survey data from 2021 to 2023, and employing Bayesian ordinal probit and Bayesian ordinal quantile modeling to evaluate the effects of these variables-while controlling for other commonly used covariates-on EV purchase intent, both on average and across its full distribution.
+By modeling purchase intent as an ordered outcome-from “not at all likely” to “very likely”-we reveal how covariate effects differ across levels of interest.
+This is the first application of ordinal quantile modeling in the EV adoption literature, uncovering heterogeneity in how potential buyers respond to key factors.
+For instance, confidence in development of charging infrastructure and belief in environmental benefits are linked not only to higher interest among likely adopters but also to reduced resistance among more skeptical respondents.
+Notably, we identify a gap between the prevalence and influence of key predictors: although few respondents report strong infrastructure confidence or frequent EV information exposure, both factors are strongly associated with increased intent across the spectrum.
+These findings suggest clear opportunities for targeted communication and outreach, alongside infrastructure investment, to support widespread EV adoption.
 keywords:
-Decarbonization, electric vehicle, ordinal probit, Pew Research, quantile regression, technology adoption.
-journal: arXiv.org
+Decarbonization; Electric vehicle; Ordinal probit; Pew Research Center; Quantile regression; Technology adoption.
+journal: Transportation Research Part A
 1 Introduction
 Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024).
-Electric vehicles (EVs) are widely regarded as a key solution for reducing emissions and meeting climate targets, given their zero tailpipe emissions.
-To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives.
-In the US, key measures include nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program..
-Despite these efforts, EV sales in the US remain relatively low, accounting for only about 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024).
-At the same time, the future trajectory of EV adoption in the US is increasingly uncertain.
+Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets (Bhat et al., 2025).
+To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives (Jenn et al., 2018; Mekky and Collins, 2024).
+Key US measures have included nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program.
+As of late 2025, several of these EV-focused policies have been rolled back under the Trump administration..
+Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024).
+The trajectory of US adoption is also increasingly uncertain.
 In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024).
-Moreover, successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations.
-The Biden administration has promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its climate agenda, while the Trump administration has signaled a deregulatory shift on EV policies with plans to weaken both incentives as well as regulations.
-[^2] President Trump’s Executive Order on January 20, 2025, outlined plans to halt \$7.5 billion in charging infrastructure funding, weaken vehicle emissions standards that drive EV production, eliminate EV tax credits, and revoke California’s EV mandate waiver, which would limit state-level regulations.
-See https://www.whitehouse.gov/presidential-actions/2025/01/unleashing-american-energy/ These shifting political priorities may reflect broader tensions between environmental regulations, industry interests, and fiscal considerations in EV adoption.
-While EVs help reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024).
-The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and restrict consumer choice.
-Ultimately, the future of the EV market depends on consumer interest, which is likely to be shaped by multiple factors, including cost considerations, convenience, perception of environmental benefits, and awareness of these trade-offs.
-Although EVs often have higher upfront costs due to expensive battery technology, they can offer long-term savings through lower operational and maintenance expenses, as electric motors require less maintenance.
-However, the overall cost of owning an EV can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024).
-Beyond cost considerations, access to reliable charging infrastructure remains an important consideration.
-Limited charging availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel.
-Environmental concerns also shape consumer attitudes, but perceptions may vary, given the regional heterogeneity in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
-Recognizing the importance of public interest in shaping the future of this market, this study examines how public opinion on purchasing EVs is associated with factors such as exposure to EV information, perceptions about the environmental impact of EVs (as compared to gas-powered vehicles) and government actions to mitigate climate change, confidence in EV related infrastructure development, prior ownership, and demographic characteristics such as age, gender, education, political affiliation, and geographical location of the respondents.
+Successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations.
+The Biden administration promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its broader climate agenda.
+In contrast, the Trump administration has pursued a deregulatory shift: as of late 2025, the federal EV tax credits are being phased out and the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked.
+[^2] For details on the EV tax credit phase-out, see Section 70502 of the One Big Beautiful Bill Act.
+For details on the status of California’s waiver for ZEV mandates, see H.J.
+Res.
+88.
+For details on the reconsideration of the endangerment finding, see the EPA website.
+Federal charging-infrastructure funding was also suspended but is slated for restoration following litigation (see Department of Transportation newsroom).
+Other policy changes are subject to ongoing or possible future legal challenges, so further changes remain possible.
+These changing political priorities reflect tensions between environmental regulations, industry interests, and fiscal considerations in promoting EV adoption.
+Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024).
+The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and limit consumer choice.
+Ultimately, the adoption of EVs depends on consumer interest, which is likely to be shaped by multiple factors, such as cost, convenience, perception of environmental benefits, and awareness of these trade-offs (Rezvani et al., 2015).
+Although EVs often have higher upfront costs due to expensive batteries, they can offer long-term savings as electric motors require less maintenance (Liu et al., 2021).
+However, the overall cost of ownership can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024).
+Access to reliable charging infrastructure also remains an important consideration (Chakraborty et al., 2019, 2022).
+Limited availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel.
+Environmental concerns also affect consumer attitudes, but perceptions may vary, given the regional differences in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
+Recognizing the importance of public interest in shaping the future of this market, we examine how stated purchase intent in the US relates to the following four key factors: exposure to EV information, perceptions of their environmental benefits (compared to gas-powered vehicles), perceptions of government action to mitigate climate change, and confidence in infrastructure development.
+At the same time, we control for prior ownership, and commonly used demographic variables such as age, gender, education, political affiliation, and geographic location.
+The data for these four focal factors, along with the dependent and control variables, are drawn from three nationally representative opinion surveys conducted by the Pew Research Center over multiple years.
+Our empirical analysis consists of two stages.
+First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how these factors relate to an ordinal measure of EV purchase intent.
+We present the covariate effects on the average probabilities of reporting each level of interest, an aspect rarely explored in the EV literature.
+Secondly, and perhaps more importantly, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how these factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined.
+The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
 Our paper makes three key contributions.
-First, we explore the role of several novel factors that have been largely overlooked in the existing EV adoption literature.
-Previous studies have examined how a combination of factors, including range anxiety, performance concerns, charging infrastructure, environmental attitudes, government policies, peer effects, and demographic characteristics, influences EV purchase intent (Carley et al., 2013; Tiwari et al., 2020; Zhao et al., 2022; Mamkhezri, 2025).
-Other studies examine the impact of specific factors, such as providing information about total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), reputation (Buhmann and Criado, 2023), and consumer incentives (Xue et al., 2023; Stekelberg and Vance, 2024).
-Related research examines factors driving actual EV sales rather than purchase intent, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023).
-We contribute to this literature by examining how EV purchase intent is influenced by factors such as confidence in future charging infrastructure development, perceptions of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits.
-Our second contribution is related to usage of large and nationally representative datasets.
-Existing research studying the trends and determinants of public opinion and purchase intent for EVs has predominantly relied on small-scale surveys.
+First, we examine four unique factors (while controlling for a host of demographic variables) to understand how informational, perceptual, and trust-related factors influence the likelihood and intensity of stated EV purchase intent in the US, and how do these effects vary across different levels of consumer interest.
+Prior studies of EV purchase intent have considered related factors, but these have either been distinct from the ones we study or investigated only in small samples.
+The factors we examine have not been analyzed jointly, on large samples, or in the US context.
+The first factor is confidence in future charging infrastructure development.
+While earlier work has emphasized the role of existing infrastructure (Carley et al., 2013; Tiwari et al., 2020), purchase intent is also likely to depend on expectations about future network expansion, particularly when the current system remains incomplete.
+The second factor is perceptions of government climate efforts.
+Past studies have examined links between purchase intent and political affiliation (e.g., Mamkhezri (2025); Sintov et al.
+(2020)), but not views on government climate action.
+Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase—especially if they are less aware of other benefits.
+The third factor is exposure to EV-related information.
+Previous research has examined the effects of information on total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), and peer effects (Zhao et al., 2022).
+We build on this by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference.
+The fourth factor is perceptions of EVs’ environmental benefits.
+Several studies have considered consumers’ pro-environmental attitudes (Hu et al., 2023; Carley et al., 2013; Krause et al., 2013; Khatua et al., 2023; Scherrer, 2023; Mamkhezri, 2025) and concern about reputation (Buhmann and Criado, 2023).
+However, research on perceptions of EV’s environmental benefits relative to gasoline powered vehicles is limited.
+Lashari et al.
+(2021) and Jia et al.
+(2025) examine this factor in a small sample, whereas our analysis situates it in a broader US context.
+[^3] Studies have also examined other factors such as range anxiety, performance concerns, and consumer incentives (See Carley et al.
+(2013); Tiwari et al.
+(2020); Zhao et al.
+(2022); Stekelberg and Vance (2024); Mamkhezri (2025) among others).
+A separate strand of research examines determinants of actual EV sales, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), peer effects (Chakraborty et al., 2022), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023).
+Our findings reveal sizable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
+Our second contribution is the use of large, nationally representative datasets, which stands in contrast to the predominantly small-scale, non-representative surveys commonly used in prior research on EV purchase intent.
 For example, Tiwari et al.
-(2020) analyzed 1,800 responses from UK residents, Lashari et al.
-(2021) analyzed 1,500 responses from South Korea, Hu et al.
-(2023) analyzed 807 responses from China, and Mamkhezri (2025) analyzed 1,500 responses from the US.
-[^3] While some recent studies (e.g., Pani et al.
-(2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of perception.
+(2020), Hu et al.
+(2023), and Mamkhezri (2025) analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the United States, respectively.
+[^4] Although some recent studies (e.g., Pani et al.
+(2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it.
 Singh et al.
-(2020) provide a meta-analysis of 211 peer-reviewed research articles published between 2009 and 2019, identifying key factors influencing public opinion of EVs.
-Notably, most such studies are either based on a single survey for a particular year, not nationally representative, or constrained by sample size; which diminishes their generalizability.
-This paper overcomes such data related limitations by drawing on three nationally representative opinion polls conducted by the Pew Research Center in the US with a significantly larger sample size than existing previous studies.
-Specifically, we use data from 2021, 2022, and 2023, with 11,052, 7,173, and 7,201 observations, respectively.
-Our third contribution lies in the application of advanced modeling and estimation techniques to study EV purchase intent.
-Existing research has used various methods to analyze EV purchase intent, such as linear regression to assess the effects of personal beliefs and preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation in shaping EV preferences (Buhmann and Criado, 2023); mixed logit models to explore attitudes towards various EV policy incentives, employment changes and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022).
-These papers rely on Classical (Frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters.
-This can lead to biased estimates of covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly translate to changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
-In contrast, this paper adopts a Bayesian modeling framework that provides a posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables us to make direct probability statements about the parameters--rather than relying on the notion of “confidence” as in the Classical framework.
-Our empirical analysis consists of two stages.
-First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how different factors are associated with consumer interest in EV adoption.
-We present the covariate effects on the average probabilities of outcomes, an aspect which is rarely explored in the EV literature.
-Next, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how different factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined.
-Notably, this study is the first to apply ordinal quantile analysis to EV purchase intent.
-Since no classical estimation methods exist for ordinal quantile models, Bayesian estimation is essential for this analysis.
-The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
+(2020) provide a meta-analysis of 211 peer-reviewed research articles 2009–2019, identifying key factors influencing public opinion of EVs.
+Notably, most such studies are based on single-year, non-representative, or small samples.
+We overcome these limitations by using three nationally representative US opinion polls, with significantly larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023.
+This allows for greater statistical power and more robust generalizability of our findings.
+Our third-and perhaps most significant-contribution lies in applying advanced modeling and estimation techniques to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent.
+Research has used various methods, such as linear regression to assess the effects of personal beliefs, preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation (Buhmann and Criado, 2023); mixed logit models to explore attitudes toward various EV policy incentives, employment changes, and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers, such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022).
+A key limitation of these approaches is that they cannot capture heterogeneity in covariate effects across the full distribution of intent.
+We address this by applying ordinal quantile analysis to EV purchase intent, an approach not previously used in this context.
+In contrast to prior work that focuses on average effects, this approach captures heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase).
+This yields more detailed evidence on how different factors shape adoption intent.
+For example, it shows that confidence in infrastructure development and belief in environmental benefits are not only associated with higher intent among likely adopters, but also with a lower probability of reporting no interest among more skeptical respondents.
+Note that no classical estimation methods exist for ordinal quantile models, so we use a Bayesian approach.
+Finally, our use of a Bayesian framework provides several advantages over traditional (frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters.
+This can bias covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly reflect changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
+In contrast, Bayesian framework allows us to obtain posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables making direct probability statements about the parameters-rather than relying on the notion of statistical confidence as in the classical framework.
 The remainder of the article is structured as follows.
-Section 2 introduces the ordinal probit and ordinal quantile models employed in our analysis, along with their MCMC algorithms for estimation.
 Section 2 outlines the data and provides a preliminary investigation.
-In Sections 4, we present the results from our models and discuss their implications.
+Section 3 introduces the ordinal probit and ordinal quantile models, along with their MCMC algorithms for estimation.
+In Section 4, we present the results from our models and in Section 5 discuss policy implications.
 Finally, Section 6 offers concluding remarks.
-2 The Modeling Framework
-In this section, we briefly describe the ordinal probit model followed by the ordinal quantile model.
-For both the models, we outline the Bayesian estimation algorithm using Markov chain Monte Carlo (MCMC) methods and explain the computation of covariate effects.
-2.1 The Ordinal Probit Model
-Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates.
-These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al.
-(2023).
-In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked.
-However, these scores only carry ordinal meaning and lack cardinal interpretation.
-Thus, while the categories are ranked, the differences between them cannot be compared directly.
-For example, in our current study, responses to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’.
-Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1.
-Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
-The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
-$z_{i}=x^{\prime}_{i}\beta+\varepsilon_{i},\hskip 54.2025pt\forall\;i=1,\cdots,n,$ (1)
-where $x_{i}$ is a $k\times 1$ covariate vector including a column of ones, $\beta$ is a $k\times 1$ parameter vector, and $n$ is the number of observations.
-We assume the error $\varepsilon$ is independent and identically distributed (iid) as a standard normal distribution i.e., $\varepsilon_{i}\sim N(0,1)$ for $i=1,2,\ldots,n$; which gives rise to the ordinal probit model.
-In our study, the unobserved variable $z_{i}$ represents intent to purchase EV and is associated with the observed discrete response $y_{i}$, that denotes response to the question about purchasing an EV, as follows:
-[…]
-where $-\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$.
-This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is visually represented in Figure 1.
-As seen from the figure, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so on.
-We set $\gamma_{1}=0$ and $var{(\varepsilon)}=1$ to anchor the location and scale of the distribution, respectively.
-These restrictions are required for parameter identification and further explained in (Jeliazkov et al., 2008) and Jeliazkov and Rahman (2012).
-Figure 1: A pictorial representation of the distribution of the latent variable $z$.
-The four probabilities P$(y_{i}=1)$, P$(y_{i}=2)$, P$(y_{i}=3)$ and P$(y_{i}=4)$ correspond to the responses: ‘not at all likely’, ‘not too likely’, ‘somewhat likely’, and ‘very likely’, regarding the intention to purchase an electric vehicle ($z_{i}$) for individual $i$ with mean $x^{\prime}_{i}\beta$.
-Given the data $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
-$\begin{split}f(y|\beta,\gamma)&=\prod_{i=1}^{n}\prod_{j=1}^{J}\Pr(y_{i}=j|\beta,\gamma)^{I(y_{i}=j)}=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}\Phi(\gamma_{j}-x^{\prime}_{i}\beta)-\Phi(\gamma_{j-1}-x^{\prime}_{i}\beta)\bigg{]}^{I(y_{i}=j)},\end{split}$ (3)
-where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise.
-But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling.
-So, we apply the transformation of Albert and Chib (2001): $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $12\leq j\leq J-$ 1, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
-To complete the Bayesian framework, we utilize the Bayes theorem to combine the augmented likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$.
-The resulting augmented posterior distribution can be written as,
-$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
-where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, the correspondence between $\gamma$ and $\delta$ is determined by the one-to-one mapping, and the cutpoint index $j$ is determined by the observed value of $y_{i}$.
-Additionally, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$ and the priors $\pi(\beta)$ and $\pi(\delta)$ are pdf’s of normal distribution.
-Based on the augmented joint posterior (4), the conditional posteriors can be derived and the model can be estimated using MCMC sampling, as outlined in Algorithm 1.
-Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effect of a variable.
-Let’s assume that the $l$-th covariate, ${x_{i,l}}$, is set to two distinct values, $a$ and $b$, denoted as ${x_{i,l}^{a}}$ and ${x_{i,l}^{b}}$, respectively.
-We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{p,l},\beta_{p,-l})$, where $-l$ denotes all covariates (and parameters) except for the $l$-th covariate (and parameter).
-Our goal is to examine the distribution of the difference ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$ for $1\leq j\leq J$, by marginalizing over ${x_{i,-l}}$ and $(\beta_{p},\delta_{p})$.
-The marginalization enables us to incorporate the uncertainty associated with data and model parameters, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
-Algorithm 1: MCMC algorithm for estimating ordinal probit model.
-  1. 1.
-    Sample $\delta,z|y,\beta$ in one block as follows:
-    1. (a)
-      Sample $\delta|y,\beta$ marginally of $z$ by generating $\delta^{\prime}$ from a random-walk chain $\delta^{\prime}=\delta+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter, $\hat{D}=\left.-\left[\frac{\partial^{2}\ln f(y|\beta,\delta)}{\partial\delta\partial\delta^{T}}\right]^{-1}\right|_{\delta=\hat{\delta}}$, and $\hat{\delta}=\arg\max_{\delta}\ln f(y|\beta,\delta)$.
-Accept the proposed value $\delta^{\prime}$ with probability
-      $\alpha_{MH}(\delta,\delta^{\prime})=\min\left\{1,\frac{f(y|\beta,\delta^{\prime})\pi(\delta^{\prime})}{f(y|\beta,\delta)\pi(\delta)}\right\},$
-      otherwise repeat the current value $\delta$.
-    2. (b)
-      Sample $z_{i}|y,\beta,\delta\sim TN_{\left(\gamma_{j-1},\gamma_{j}\right)}\left(x_{i}^{\prime}\beta,1\right)$ for $i=1,\ldots,n$, where $TN_{(\gamma_{j-1},\gamma_{j})}$ denotes a truncated normal distribution constrained between $\gamma_{j-1}$ and $\gamma_{j}$; and $\gamma$ is derived from the one-to-one mapping relating $\gamma$ and $\delta$.
-  2. 2.
-    Sample $\beta|z\sim N(\hat{\beta},\hat{B})$ where $\hat{B}^{-1}=({B}_{0}^{-1}+X^{\prime}X)$, and $\hat{\beta}=\hat{B}(B_{0}^{-1}\beta_{0}+X^{\prime}z)$.
-To obtain draws from the distribution ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$, we use the method of composition (Chib and Jeliazkov, 2006).
-This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference ${\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)}$, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$.
-The process is repeated for all individuals and MCMC draws.
-The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
-$ACE_{y=j}\approx\frac{1}{M}\frac{1}{n}\sum_{m=1}^{M}\sum_{i=1}^{n}\Big{[}\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta^{(m)},\delta^{(m)})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta^{(m)},\delta^{(m)})\Big{]},$ (5)
-where $(\beta^{(m)},\delta^{(m)})$ denotes the MCMC draws, and $M$ is the number of post burn-in MCMC draws.
-2.2 The Ordinal Quantile Model
-The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates.
-However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions.
-For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs.
-However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers.
-To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression.
-These insights provide suggestive evidence for more targeted interventions; for example, they show that improving infrastructure visibility may be more effective when directed at consumers already considering EVs, rather than those with little or no interest.
-The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
-$z_{i}$ $=x^{\prime}_{i}\beta_{p}+\varepsilon_{i},$ $\qquad\forall\;i=1,\cdots,n,$ (6)
-[…]
-where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005).
-The AL distribution is essential to create a working likelihood function, and yields the following expression,
-$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg{]}^{I(y_{i}=j)},\end{split}$ (7)
-where $F_{AL}(\cdot)$ denotes the cdf of an AL distribution and $I(y_{i}=j)$ is an indicator function.
-At this stage, three key issues need attention.
-First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $12\leq j\leq J-$ 1, where note that $\delta_{p}$ is quantile dependent.
-Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is fixed for any quantile to meet the scale restriction.
-Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors.
-So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$.
-The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$.
-Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
-By Bayes theorem, we combine the augmented likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at the augmented joint posterior distribution,
-$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big{\{}\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big{\}}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg{\{}\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg{\}}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
-In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$.
-In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1.
-In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
-The conditional posterior distributions are derived from the augmented joint posterior distribution (i.e.
-equation 8), and the parameters are sampled as per Algorithm 2.
-The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
-Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
-  1. 1.
-    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
-    $\tilde{B}^{-1}_{p}=\bigg{(}\sum_{i=1}^{n}\frac{x_{i}x^{\prime}_{i}}{\tau^{2}w_{i}}+B_{p0}^{-1}\bigg{)},\quad\mbox{and}\quad\tilde{\beta}_{p}=\tilde{B}_{p}\bigg{(}\sum_{i=1}^{n}\frac{x_{i}(z_{i}-\theta w_{i})}{\tau^{2}w_{i}}+B_{p0}^{-1}\beta_{p0}\bigg{)}.$
-  2. 2.
-    Sample $w_{i}|\beta_{p},z_{i}$ $\sim$ $GIG\,(0.5,\tilde{\lambda}_{i},\tilde{\eta})$, for $i=1,\cdots,n$, where,
-    $\tilde{\lambda}_{i}=\Big{(}\frac{z_{i}-x^{\prime}_{i}\beta_{p}}{\tau}\Big{)}^{2},\quad\mbox{and}\quad\tilde{\eta}=\Big{(}\frac{\theta^{2}}{\tau^{2}}+2\Big{)}.$
-  3. 3.
-    Sample $\delta_{p}|y,\beta_{p}$ marginally of $w$ (latent weight) and $z$ (latent data), by generating $\delta_{p}^{\prime}$ from a random-walk chain $\delta^{\prime}_{p}=\delta_{p}+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter and $\hat{D}$ denotes negative inverse Hessian, obtained by maximizing the log-likelihood with respect to $\delta_{p}$.
-Given the current value of $\delta_{p}$, accept the proposed draw $\delta^{\prime}_{p}$ with probability,
-    $\alpha_{MH}(\delta_{p},\delta^{\prime}_{p})=\min\bigg{\{}1,\frac{\;f(y|\beta_{p},\delta^{\prime}_{p})\;\pi(\beta_{p},\delta^{\prime}_{p})}{f(y|\beta_{p},\delta_{p})\;\pi(\beta_{p},\delta_{p})}\bigg{\}};$
-    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
-  4. 4.
-    Sample $z_{i}|y,\beta_{p},\gamma_{p},w$ $\sim$ $TN_{(\gamma_{p,j-1},\gamma_{p,j})}(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\cdots,n$, where $TN$ denotes a truncated normal distribution and $\gamma_{p}$ is obtained via $\delta_{p}$ using the one-to-one mapping.
-After estimating the ordinal quantile models, we compute the covariate effects using the method of composition.
-Existing studies on quantile regression that have applied this method to estimate the covariate effects include (Rahman and Vossmeyer, 2019) and Bresson et al.
-(2021).
-Assuming a setup similar to Section 2.1, our goal is to obtain draws from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$.
-We randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$.
-This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
 2 Data
 The study draws on survey data from three opinion polls carried out by the Pew Research Center over a three-year span.
-The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second survey took place during May 2-8, 2022, on a sample of 10,282 adults, and the third survey was administered during May 30-June 4, 2023 on a sample of 10,329 adults.
-After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis.
+The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second during May 2-8, 2022 on a sample of 10,282 adults, and the third during May 30-June 4, 2023 on a sample of 10,329 adults.
+After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis [^5] All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage)..
 Table 1 shows a descriptive summary of the independent and dependent variables.
 Specifically, it presents the mean and standard deviation of the continuous variables, and count and percentage for the categorical variables.
-The above survey data offer at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent.
-First, our survey data is nationally representative and span multiple years.
+The data utilized offers at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent.
+First, they are nationally representative and span multiple years.
 As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity.
-All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage).
-Second, an appealing feature of our data sets is the availability of ordinal responses to likelihood of purchasing EV.
-This enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023).
-Third, we note that the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase.
+Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023).
+Third, the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase.
 Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
@@
 count per count per count per
-EV Info $\circ$ Read or heard a lot 4170 37.73 ..absent... . ..absent... . ..absent... . ..absent... .
-$\circ$ Little or nothing at all 6882 62.27 ..absent... . ..absent... . ..absent... . ..absent... .
-Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 ..absent... . ..absent... . ..absent... . ..absent... .
-$\circ$ About the same or worse 3108 28.12 ..absent... . ..absent... . ..absent... . ..absent... .
-EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident ..absent... . ..absent... . ..absent... . ..absent... . 1263 17.54
-$\circ$ Somewhat confident, not too confident, or not at all confident ..absent... . ..absent... . ..absent... . ..absent... . 5938 82.46
+EV Info $\circ$ Read or heard a lot 4170 37.73 .... .... .... ....
+$\circ$ Little or nothing at all 6882 62.27 .... .... .... ....
+Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 .... .... .... ....
+$\circ$ About the same or worse 3108 28.12 .... .... .... ....
+EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident .... .... .... .... 1263 17.54
+$\circ$ Somewhat confident, not too confident, or not at all confident .... .... .... .... 5938 82.46
 EV Owner Owns electric or hybrid vehicle 1039 9.40 769 10.72 812 11.28
@@
 $\circ$ Too little 6887 62.31 3853 53.72 3883 53.92
-Age $<50$ Indicator for age below 50 5124 46.36 2687 37.46 2596 36.05
+Age <50<50 Indicator for age below 50 5124 46.36 2687 37.46 2596 36.05
 Female Indicator for female gender 5848 52.91 3819 53.24 3815 52.98
@@
 $\circ$ HS and below 1411 12.77 1126 15.70 1195 16.59
-Married Married or living with a partner ..absent... . ..absent... . 5692 79.35 5640 78.32
+Married Married or living with a partner .... .... 5692 79.35 5640 78.32
 Metropolitan Lives in a metropolitan area 9817 88.83 6260 87.87 6286 87.29
@@
 Not at all likely 2441 22.09 1924 26.82 2253 31.29
-We begin by discussing the independent variables that have a direct influence on purchasing an EV–the question of primary interest in this article.
+We begin by discussing the four independent variables that are of primary interest in this article.
 Two such variables, ‘EV Info’ and ‘Env Better’, are specific to the April 2021 survey.
@@
 According to Table 1, approximately 72% of respondents view EVs as being environmentally superior to gas-powered vehicles.
-Another noteworthy variable is ‘EV Infra’, which is only available in the May-June 2023 survey.
+The third variable of interest ‘EV Infra’ is only available in the May-June 2023 survey.
 This indicator variable is coded as 1 if the respondent is extremely or very confident that the US will build the necessary charging stations and infrastructure to support large number of EVs on the roads, and 0 otherwise (somewhat confident, not too confident, or not at all confident).
 As seen in Table 1, about 82% of respondents express a negative perception towards the government for creating the infrastructure needed to support the adoption of EVs.
-Two key variables of interest to our study, which are available in all the surveys, are ‘EV Owner’ and ‘GACC’.
-The ‘EV Owner’ variable indicates whether the respondent currently owns an electric or hybrid vehicle and takes the value 1 if true and 0 otherwise.
-As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets.
-The ‘GACC’ variable reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC).
+The fourth variable of interest ‘GACC’ reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC) and is available in all three surveys.
 This variable is divided into three categories: ‘too much’, ‘about the right amount’, and ‘too little’ (with ‘too little’ serving as the reference category).
 Table 1 indicates that in each sample, more than half of the respondents feel that government efforts to combat global climate change are insufficient.
+Another variable, which is an important predictor for EV purchase, is EV Owner, indicating whether the respondent currently owns an electric or hybrid vehicle and taking the value 1 if true and 0 otherwise.
+As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets.
 We also incorporate a broad set of demographic variables as covariates in our regression models.
@@
 We now examine the construction of these variables in more detail.
-Pew surveys record income as belonging to one of nine categories: $0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k$, and $>100k$, where $k$ denotes a thousand dollars.
+Pew surveys record income as belonging to one of nine categories: 030k,30k40k,40k50k,50k60k,60k70k,70k80k,80k90k,90k100k0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k, and >100k>100k, where $k$ denotes a thousand dollars.
 To simplify, we use the mid-point of each income range and assign an imputed value of \$20,000 for the first category and \$110,000 for the last category.
@@
 Therefore, respondents across the three samples are generally well-educated.
-Information on marital status is available only in the May 2022 and May--June 2023 surveys.
+Information on marital status is available only in the May 2022 and May-June 2023 surveys.
 We create an indicator variable ‘Married’ that takes value 1 if the respondent is married or living with a partner, and 0 if they are divorced, separated, widowed, or never been married.
@@
 Respondents are roughly equally distributed across these three political categories.
+Figure 1: Stacked bars displaying the percentage of observations corresponding to the four opinion categories for a selected list of covariates.
+In each panel, the first stacked bar shows the percentage of observations for ‘not at all likely’ and the cumulative percentage for ‘not at all likely’ and ‘not too likely’.
+The second stacked bar shows the percentage of observations for ‘somewhat likely’ and the cumulative percentage for ‘somewhat likely’ and ‘very likely’.
 As noted earlier, our objective is to analyze the factors that influence individuals’ decision to purchase an EV.
@@
 Overall, our data summary clearly shows that public attitude towards purchasing an EV is generally positive.
-Figure 2: Stacked bars displaying the percentage of observations corresponding to the four opinion categories for a selected list of covariates.
-In each panel, the first stacked bar shows the percentage of observations for ‘not at all likely’ and the cumulative percentage for ‘not at all likely’ and ‘not too likely’.
-The second stacked bar shows the percentage of observations for ‘somewhat likely’ and the cumulative percentage for ‘somewhat likely’ and ‘very likely’.
 Further insights into the association between the dependent variable and selected regressors from the three survey datasets can be gained from Figure 1.
@@
 With respect to gender, the likelihood of purchasing an EV among females is 53.27% in the April 2021 dataset, decreases to 49.18% in May 2022, and further drops to 44.69% in the May-June 2023 dataset.
-4 Empirical Results and Discussion
-In this section, we report the results from the Bayesian estimation of ordinal probit and ordinal quantile models.
-We also report the marginal likelihood which may be used for model comparison.
-Morevoer, for each model, we compute the covariate effects for all or some of the significant variables and discuss our findings in the context of existing literature.
+3 The Modeling Framework
+In this section, we briefly describe the ordinal probit model followed by the ordinal quantile model.
+For both the models, we outline the Bayesian estimation algorithm using Markov chain Monte Carlo (MCMC) methods and explain the computation of covariate effects.
+3.1 The Ordinal Probit Model
+Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates.
+These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al.
+(2023).
+In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked, without any cardinal interpretation.
+For example, in our study, responses ($y$) to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’.
+Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1.
+Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
+The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
+$z_{i}=x^{\prime}_{i}\beta+\varepsilon_{i},\hskip 54.2025pt\forall\;i=1,\cdots,n,$ (1)
+where $x_{i}$ is a $k\times 1$ covariate vector including a column of ones, $\beta$ is a $k\times 1$ parameter vector, and $n$ is the number of observations.
+We assume the error $\varepsilon$ is independent and identically distributed (iid) as a standard normal distribution i.e., $\varepsilon_{i}\sim N(0,1)$ for $i=1,2,\ldots,n$; which gives rise to the ordinal probit model.
+In our study, the unobserved variable $z_{i}$ represents an individual’s intent to purchase EV and is linked to the observed ordinal response $y_{i}$ as follows:
+$\gamma_{j-1}<z_{i}\leq\gamma_{j}\;\longleftrightarrow\;\emph{$y_{i}$ = j},\hskip 54.2025pt\forall\;i=1,\cdots,n;\;j=1,\cdots,J,$ (2)
+where $\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$.
+This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is illustrated in Figure 2.
+As shown, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so forth.
+To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively (Jeliazkov et al., 2008; Jeliazkov and Rahman, 2012).
+[…]
+Given the responses $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
+$\begin{split}f(y|\beta,\gamma)&=\prod_{i=1}^{n}\prod_{j=1}^{J}\Pr(y_{i}=j|\beta,\gamma)^{I(y_{i}=j)}=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[\Phi(\gamma_{j}-x^{\prime}_{i}\beta)-\Phi(\gamma_{j-1}-x^{\prime}_{i}\beta)\bigg]^{I(y_{i}=j)},\end{split}$ (3)
+where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise.
+But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling.
+So, following Albert and Chib (2001), we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
+To complete the Bayesian framework, we utilize the Bayes theorem to combine the complete data likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$.
+The resulting complete data posterior can be written as,
+$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
+where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors.
+Given the complete data posterior (4), the conditional posteriors can be derived and the model estimated using MCMC sampling as outlined in Algorithm 1.
+Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effects.
+Let the $l$-th covariate, $x_{i,l}$}, take two distinct values, $a$ and $b$, denoted as $x_{i,l}^{a}$} and $x_{i,l}^{b}$}, respectively.
+We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{l},\beta_{-l})$, where l-l denotes all covariates (and parameters) except for the $l$-th covariate (and parameter).
+We evaluate the distribution of the difference $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$} for $1\leq j\leq J$, by marginalizing over $x_{i,-l}$} and $(\beta,\delta)$.
+This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
+Algorithm 1: MCMC algorithm for estimating ordinal probit model.
+  1. 1.
+    Sample $\delta,z|y,\beta$ in one block as follows:
+    1. (a)
+[…]
+      $\alpha_{MH}(\delta,\delta^{\prime})=\min\left\{1,\frac{f(y|\beta,\delta^{\prime})\pi(\delta^{\prime})}{f(y|\beta,\delta)\pi(\delta)}\right\},$
+      otherwise repeat the current value $\delta$.
+    2. (b)
+      Sample $z_{i}|y,\beta,\delta\sim TN_{\left(\gamma_{j-1},\gamma_{j}\right)}\left(x_{i}^{\prime}\beta,1\right)$ for $i=1,\ldots,n$, where $TN_{(\gamma_{j-1},\gamma_{j})}$ denotes a truncated normal distribution constrained between $\gamma_{j-1}$ and $\gamma_{j}$; and $\gamma$ is derived from the one-to-one mapping relating $\gamma$ and $\delta$.
+  2. 2.
+[…]
+To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition (Jeliazkov et al., 2008).
+This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference $\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)$}, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$.
+The process is repeated for all individuals and MCMC draws.
+The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
+$ACE_{y=j}\approx\frac{1}{M}\frac{1}{n}\sum_{m=1}^{M}\sum_{i=1}^{n}\Big[\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta^{(m)},\delta^{(m)})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta^{(m)},\delta^{(m)})\Big],$ (5)
+where $(\beta^{(m)},\delta^{(m)})$ denotes the MCMC draws, and $M$ is the number of post burn-in MCMC draws.
+3.2 The Ordinal Quantile Model
+The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates.
+However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions.
+For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs.
+However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers.
+To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression.
+The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
+$z_{i}$ $=x^{\prime}_{i}\beta_{p}+\varepsilon_{i},$ $\qquad\forall\;i=1,\cdots,n,$ (6)
+$\gamma_{p,j-1}$ $<z_{i}\leq\gamma_{p,j}\;\longleftrightarrow\;\emph{$y_{i}$ = j},$ $\qquad\forall\;i=1,\cdots,n;\;j=1,\cdots,J,$
+where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005).
+The AL distribution is essential to create a working likelihood function, and yields the following expression,
+$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg]^{I(y_{i}=j)},\end{split}$ (7)
+where $F_{AL}(\cdot)$ denotes the cdf of an AL distribution and $I(y_{i}=j)$ is an indicator function.
+At this stage, three key issues need attention.
+First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $2\leq j\leq J-1$.
+Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is automatically fixed for any quantile to meet the scale restriction.
+Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors.
+So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$.
+The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$.
+Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
+By Bayes theorem, we combine the complete data likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at,
+$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big\{\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big\}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg\{\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg\}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
+In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$.
+In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1.
+In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
+Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
+  1. 1.
+    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
+    $\tilde{B}^{-1}_{p}=\bigg(\sum_{i=1}^{n}\frac{x_{i}x^{\prime}_{i}}{\tau^{2}w_{i}}+B_{p0}^{-1}\bigg),\quad\mbox{and}\quad\tilde{\beta}_{p}=\tilde{B}_{p}\bigg(\sum_{i=1}^{n}\frac{x_{i}(z_{i}-\theta w_{i})}{\tau^{2}w_{i}}+B_{p0}^{-1}\beta_{p0}\bigg).$
+  2. 2.
+[…]
+    $\tilde{\lambda}_{i}=\Big(\frac{z_{i}-x^{\prime}_{i}\beta_{p}}{\tau}\Big)^{2},\quad\mbox{and}\quad\tilde{\eta}=\Big(\frac{\theta^{2}}{\tau^{2}}+2\Big).$
+  3. 3.
+[…]
+    $\alpha_{MH}(\delta_{p},\delta^{\prime}_{p})=\min\bigg\{1,\frac{\;f(y|\beta_{p},\delta^{\prime}_{p})\;\pi(\beta_{p},\delta^{\prime}_{p})}{f(y|\beta_{p},\delta_{p})\;\pi(\beta_{p},\delta_{p})}\bigg\};$
+    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
+  4. 4.
+    Sample $z_{i}|y,\beta_{p},\gamma_{p},w$ $\sim$ $TN_{(\gamma_{p,j-1},\gamma_{p,j})}(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\cdots,n$, where $TN$ denotes a truncated normal distribution and $\gamma_{p}$ is obtained via $\delta_{p}$ using the one-to-one mapping.
+The conditional posterior distributions are derived from the complete data posterior (i.e., equation 8) and the parameters are sampled as per Algorithm 2.
+The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
+After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in (Rahman and Vossmeyer, 2019) and Bresson et al.
+(2021).
+Assuming a setup similar to Section 3.1, our goal is to draw from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$.
+So, we randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$.
+This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
+4 Empirical Results
 4.1 Results from Ordinal Probit Models
-We begin by presenting the results from ordinal probit models estimated separately for each survey data.
-Specifically, Table 2 presents the posterior means and standard deviations of the parameters along with the logarithm of marginal likelihood (Chib and Jeliazkov, 2001).
-The estimates are obtained from 10,000 MCMC iterations (after a burn-in of 2,500 iterations) and utilizes the following prior distributions: $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\ast I)$, where $I$ denotes an identity covariance matrix.
-The priors are relatively diffuse to allow the data speak for itself.
-In ordinal models, the sign of coefficients indicates the direction of covariate effects for the first and last response categories.
-For example, in Table 2, the coefficient for “EV Info” is positive (0.43), which implies that it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”) to purchase an EV.
-However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the covariate effects on intermediate outcomes (i.e., “Not too likely” and “Somewhat likely”).
-Therefore, Table 3 provides the covariate effects for all variables which are statistically different from zero.
-Note, our measure of covariate effects account for uncertainty in parameters and observations to avoid biases (Jeliazkov and Vossmeyer, 2018).
-The combined results from the 2021, 2022, and 2023 surveys reveal consistent patterns in factors associated with EV purchase intent.
-As shown in Table 3, the covariate effects corresponding to the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs as compared to those with limited exposure to EV-related information.
+We begin by presenting the results from ordinal probit models estimated separately for each survey.
+Table 2 presents the posterior means and standard deviations of the parameters based on 10,000 MCMC iterations after a burn-in of 2,500 iterations.
+We utilize the priors $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\times I)$, where $I$ denotes an identity covariance matrix.
+These priors are relatively diffuse, allowing the data to primarily drive the posterior inference.
+Table 2 also reports the logarithm of marginal likelihood, which is the most accepted quantity for Bayesian model comparison (Chib and Jeliazkov, 2001; Jeliazkov et al., 2008).
+The marginal likelihood (also known as model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution.
+It accounts for both model fit and complexity, with higher values indicating a better model fit.
+The use of marginal likelihood contrasts with those commonly used in the classical framework, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) (Lin and Tan, 2017; Kumar et al., 2025).
+In ordinal models, coefficient signs indicate the direction of covariate effects for the first and last response categories.
+For example, in Table 2, ‘EV Info’ has a positive coefficient (0.43), so it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”).
+However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the effects on intermediate outcomes (“Not too likely” and “Somewhat likely”).
+Therefore, Table 3 provides the covariate effects for all variables whose parameters are statistically different from zero at 95% probability level.
+[^6] As noted earlier, some variables of interest are available in only one year.
+For variables measured in all three survey years, however, the estimated effects are similar across years, suggesting that the influences on EV purchase intent are relatively stable over time.
+This stability supports the credibility of estimates for variables observed in a single year.
+The combined results from the three surveys reveal consistent patterns in factors associated with EV purchase intent.
+As shown in Table 3, the covariate effects from the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs, compared to those with limited exposure.
 These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories.
-Our findings align with Zhao et al.
-(2022) and Scherrer (2023), who demonstrate that greater exposure to information about EVs—whether through peer effects or media—positively influences attitudes toward EVs, and with Krause et al.
-(2013), who underscore the crucial role of accurate information in shaping consumer perceptions.
+Our findings complements Zhao et al.
+(2022) and Scherrer (2023), who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al.
+(2013), who underscore the role of accurate information in shaping perceptions.
 Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability.
-For the 2021 survey, we also observe, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent.
-This finding aligns with Carley et al.
+For the 2021 survey, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent.
+This finding complements Carley et al.
 (2013), Krause et al.
-(2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs  [^4] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
-Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood, from April 2021; May 2022; and May–June 2023 survey data.
+(2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs.
+Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^7] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
+Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood ($\ln$ ML), from April 2021; May 2022; and May–June 2023 survey data.
 April 2021 May 2022 May–June 2023
 Mean Std Mean Std Mean Std
-Intercept $-0.28$ $0.07$ $0.36$ $0.09$ $-0.05$ $0.09$
-EV Info $0.43$ $0.02$.absent... . ..absent... . ..absent... . ..absent... .
-Env Better $0.79$ $0.03$.absent... . ..absent... . ..absent... . ..absent... .
-EV Infra ..absent... . ..absent... . ..absent... . ..absent... . $0.82$ $0.04$
+Intercept 0.28-0.28 $0.07$ $0.36$ $0.09$ 0.05-0.05 $0.09$
+EV Info $0.43$ $0.02$ .... .... .... ....
+Env Better $0.79$ $0.03$ .... .... .... ....
+EV Infra .... .... .... .... $0.82$ $0.04$
+GACC (TM) 0.79-0.79 $0.04$ 1.13-1.13 $0.04$ 1.07-1.07 $0.04$
+GACC (RA) 0.52-0.52 $0.03$ 0.63-0.63 $0.03$ 0.49-0.49 $0.04$
 EV Owner $0.79$ $0.04$ $0.86$ $0.05$ $0.82$ $0.04$
-GACC (TM) $-0.79$ $0.04$ $-1.13$ $0.04$ $-1.07$ $0.04$
-GACC (RA) $-0.52$ $0.03$ $-0.63$ $0.03$ $-0.49$ $0.04$
-Age $<50$ $0.16$ $0.02$ $0.17$ $0.03$ $0.15$ $0.03$
-Female $-0.02$ $0.02$ $-0.11$ $0.03$ $-0.12$ $0.03$
+Age <50<50 $0.16$ $0.02$ $0.17$ $0.03$ $0.15$ $0.03$
+Female 0.02-0.02 $0.02$ 0.11-0.11 $0.03$ 0.12-0.12 $0.03$
 Grad & Above $0.38$ $0.04$ $0.45$ $0.04$ $0.49$ $0.04$
 Some College $0.18$ $0.04$ $0.23$ $0.04$ $0.21$ $0.04$
-Married ..absent... . ..absent... . $0.06$ $0.03$ $0.02$ $0.04$
+Married .... .... $0.06$ $0.03$ $0.02$ $0.04$
 Metropolitan $0.24$ $0.03$ $0.25$ $0.04$ $0.20$ $0.04$
-Midwest $-0.08$ $0.03$ $-0.05$ $0.04$ $-0.07$ $0.04$
-South $-0.05$ $0.03$ $-0.02$ $0.04$ $-0.07$ $0.04$
+Midwest 0.08-0.08 $0.03$ 0.05-0.05 $0.04$ 0.07-0.07 $0.04$
+South 0.05-0.05 $0.03$ 0.02-0.02 $0.04$ 0.07-0.07 $0.04$
 West $0.19$ $0.03$ $0.19$ $0.04$ $0.18$ $0.04$
-US Born $-0.21$ $0.04$ $-0.33$ $0.04$ $-0.22$ $0.04$
-Black $-0.28$ $0.04$ $-0.14$ $0.06$ $-0.18$ $0.05$
-Other Races $0.05$ $0.03$ $0.01$ $0.04$ $-0.03$ $0.04$
+US Born 0.21-0.21 $0.04$ 0.33-0.33 $0.04$ 0.22-0.22 $0.04$
+Black 0.28-0.28 $0.04$ 0.14-0.14 $0.06$ 0.18-0.18 $0.05$
+Other Races $0.05$ $0.03$ $0.01$ $0.04$ 0.03-0.03 $0.04$
 Democrat $0.59$ $0.03$ $0.65$ $0.04$ $0.75$ $0.04$
@@
 Income/10000 (USD) $0.01$ $0.01$ $0.02$ $0.01$ $0.03$ $0.01$
-$\delta_{1}$ $0.01$ $0.02$ $-0.10$ $0.02$ $-0.11$ $0.02$
-$\delta_{2}$ $0.12$ $0.01$ $-0.00$ $0.02$ $0.04$ $0.02$
-$\ln$ ML $-12415$ $-8301$ $-7914$
+$\delta_{1}$ $0.01$ $0.02$ 0.10-0.10 $0.02$ 0.11-0.11 $0.02$
+$\delta_{2}$ $0.12$ $0.01$ 0.00-0.00 $0.02$ $0.04$ $0.02$
+$\ln$ ML 12415-12415 8301-8301 7914-7914
 Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest.
-Respondents who are very or extremely confident that the US will build the necessary charging infrastructure are 19.1 percentage points more likely to express strong interest in purchasing EVs and 18.8 percentage points less likely to report no interest compared to those who report lower confidence.
-This finding aligns with prior research, such as Carley et al.
+Respondents very or extremely confident that the US will build the necessary infrastructure are 19.1 (18.8) percentage points more likely to express strong (no) interest in purchasing EVs, compared to those who report lower confidence.
+This finding complements research, such as Carley et al.
 (2013) and Li et al.
-(2017).
-As shown in Table 1, 82% of respondents lack confidence that US will develop the infrastructure needed to support EV adoption.
-Taken together, these results suggest that skepticism about charging infrastructure is likely to be a significant barrier to broader adoption, highlighting the importance of both improving infrastructure and clearly communicating its availability, reliability, and expansion plans to the public.
-Table 3: Covariate effects (CE) from ordinal probit models. CE for variables not statistically different from zero are marked with $\times$.
-April 2021 May 2022 May–June 2023
-$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
-EV Info $-0.0851$ $-0.0388$ $0.0217$ $0.1022$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
-Env Better $-0.1955$ $-0.0565$ $0.0970$ $0.1551$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
-EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $-0.1880$ $-0.0553$ $0.0526$ $0.1908$
-EV Owner $-0.1321$ $-0.0849$ $0.0051$ $0.2119$ $-0.1749$ $-0.0794$ $0.0231$ $0.2312$ $-0.1814$ $-0.0513$ $0.0421$ $0.1905$
-GACC (TM) $0.1971$ $0.0453$ $-0.0945$ $-0.1478$ $0.3341$ $0.0133$ $-0.1507$ $-0.1968$ $0.3170$ $-0.0153$ $-0.1395$ $-0.1622$
-GACC (RA) $0.1180$ $0.0391$ $-0.0490$ $-0.1082$ $0.1649$ $0.0267$ $-0.0639$ $-0.1278$ $0.1282$ $0.0079$ $-0.0486$ $-0.0875$
-Age $<50$ $-0.0326$ $-0.0130$ $0.0099$ $0.0356$ $-0.0410$ $-0.0090$ $0.0129$ $0.0371$ $-0.0371$ $-0.0036$ $0.0126$ $0.0282$
-Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ $-0.0082$ $-0.0240$ $0.0301$ $0.0027$ $-0.0100$ $-0.0228$
-Grad & Above $-0.0815$ $-0.0322$ $0.0276$ $0.0860$ $-0.1139$ $-0.0240$ $0.0378$ $0.1001$ $-0.1264$ $-0.0130$ $0.0445$ $0.0949$
-Some College $-0.0373$ $-0.0152$ $0.0099$ $0.0426$ $-0.0541$ $-0.0118$ $0.0156$ $0.0503$ $-0.0517$ $-0.0052$ $-0.0166$ $0.0403$
-Married ..absent... . ..absent... . ..absent... . ..absent... . $-0.0154$ $-0.0029$ $0.0050$ $0.0134$$\times$ $\times$ $\times$ $\times$
-Metropolitan $-0.0534$ $-0.0179$ $0.0191$ $0.0522$ $-0.0662$ $-0.0099$ $0.0232$ $0.0528$ $-0.0512$ $-0.0027$ $0.0183$ $0.0356$
-Midwest $0.0162$ $0.0061$ $-0.0051$ $-0.0172$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-South $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-West $-0.0378$ $-0.0164$ $0.0101$ $0.0441$ $-0.0457$ $-0.0109$ $0.0135$ $0.0431$ $-0.0454$ $-0.0053$ $0.0148$ $0.0360$
-US Born $0.0416$ $0.0186$ $-0.0106$ $-0.0497$ $0.0777$ $0.0217$ $-0.0214$ $-0.0781$ $0.0551$ $0.0071$ $-0.0178$ $-0.0444$
-Black $0.0614$ $0.0200$ $-0.0220$ $-0.0594$ $0.0349$ $0.0058$ $-0.0116$ $-0.0291$ $0.0460$ $0.0026$ $-0.0160$ $-0.0326$
-Other Races $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Democrat $-0.1211$ $-0.0578$ $0.0403$ $0.1386$ $-0.1569$ $-0.0497$ $0.0513$ $0.1553$ $-0.1882$ $-0.0353$ $0.0677$ $0.1558$
-Ind & Others $-0.0808$ $-0.0310$ $0.0219$ $0.0898$ $-0.0890$ $-0.0187$ $0.0266$ $0.0811$ $-0.1091$ $-0.0093$ $0.0356$ $0.0827$
-$\Delta$I=\$$=\$50,000$$\times$ $\times$ $\times$ $\times$ $-0.0223$ $-0.0051$ $0.0066$ $0.0209$ $-0.0363$ $-0.0045$ $0.0116$ $0.0291$
-Across all three surveys, previous EV ownership is consistently linked with a higher interest in purchasing an EV.
-Respondents who have owned an EV or hybrid vehicle are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest compared to those without prior ownership.
-Other studies such as Carley et al.
-(2013) and Buhmann and Criado (2023) have also documented this pattern, suggesting that firsthand experience reduces uncertainty and fosters positive perceptions of EVs.
-However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in the technology, making them more engaged with the EV market.
+(2017), which highlight the importance of existing infrastructure.
 Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest.
-Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though the effects are smaller.
-Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs compared to older respondents.
+Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though to a lesser extent.
+This relationship has not, to our knowledge, been examined in prior work.
+Across all surveys, previous EV ownership is consistently linked with a higher purchase interest.
+Respondents who have owned an EV or hybrid are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest than those without prior ownership.
+This pattern, also noted by Carley et al.
+(2013) and Buhmann and Criado (2023), suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs.
+However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in EVs.
+Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs than older respondents.
 Our finding aligns with Zhao et al.
@@
 For instance, Zhao et al.
-(2022) finds that EV buyers in Shanghai are predominantly younger, with most between the ages of 31 and 40.
-Results from the 2022 and 2023 surveys also show that female respondents are modestly less likely to express strong interest and more likely to report no interest than males.
-The covariate effects on strong interest decrease between 2.3–2.4 percentage points; while the covariate effect for no interest increase between 2.7–3.0 percentage points.
-This finding is consistent several studies, including Carley et al.
+(2022) finds that EV buyers in Shanghai are mostly aged 31-40.
+Results from the 2022 and 2023 surveys also show that female respondents are are 2.3–2.4 percentage points less likely to express strong interest, and 2.7–3.0 percentage points more likely to report no interest, than males.
+This finding is consistent with Carley et al.
 (2013), Krause et al.
-(2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men in their samples exhibit a stronger preference for EVs.
+(2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men show stronger preference for EVs.
 Higher education is also linked to greater EV interest.
-Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest compared to those with a high school diploma or less.
+Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest than those with a high school diploma or less.
 Similar, though smaller, effects are observed for respondents with some college education.
-The positive relationship between education and EV interest is similarly documented in Carley et al.
+Such a relationship is similarly documented in Carley et al.
 (2013), Krause et al.
@@
 Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest.
-This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report EV preference to be higher among urban residents.
-Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest compared to those in the Northeast, though those in the Midwest and South show no significant difference from their Northeastern counterparts.
-This pattern likely reflects California’s leadership in EV adoption, driven in part by policies such as ZEV mandates and the Clean Vehicle Rebate Program, along with a robust charging network.
+This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report higher EV preference among urban residents.
+Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest than those in the Northeast; those in the Midwest and South show no significant difference from Northeastern respondents.
+This pattern likely reflects California’s leadership in EV adoption, driven in part ZEV mandates, the Clean Vehicle Rebate Program, and a robust charging network.
 [^8] See https://www.eia.gov/todayinenergy/detail.php?id=61082
-US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest compared to non-US-born respondents, possibly reflecting differences in cultural attitudes, driving patterns, or perceptions about the cost of vehicle ownership.
-Such a difference in perception about EVs is also evident in the racial background of respondents.
-Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents, likely reflecting structural barriers, such as limited access to charging infrastructure (Hsu and Fingerman, 2021).
+US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest than non-US-born respondents.
+Moreover, Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents.
 Other non-white respondents show similar preferences to White respondents.
+These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access (Hsu and Fingerman, 2021).
 Political affiliation is a consistent and significant predictor of EV purchase intent.
-Democrats are 13.9–18.8 percentage points more likely to report strong interest in EVs and 12.1–18.8 percentage points less likely to report no interest compared to Republicans.
-Similarly, Independents and others are also about 8 percentage points more likely to express strong interest in purchasing EVs relative to Republicans.
-These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible compared to Republicans, as documented by Sintov et al.
+Democrats are 13.9–18.8 percentage points more likely to report strong interest in purchasing EVs and 12.1–18.8 percentage points less likely to report no interest than Republicans.
+Independents and others are about 8 percentage points more likely to express strong interest than Republicans.
+These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible, as documented by Sintov et al.
 (2020).
-They also reflect the starkly contrasting policy approaches of the two major political parties, particularly during the tenures of Presidents Biden and Trump.
-During his first term (2017–2021), President Trump’s administration rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs.
-In contrast, the Biden administration adopted an aggressive pro-EV stance as part of its broader climate agenda, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
-Conditional on other factors, an additional \$50,000 in income is either not significantly associated with EV interest or changes the probability of strong interest (no interest) by a meagre 2.1–2.9 (2.2–3.6) percentage points.
+They also reflect the stark policy differences between the two major political parties, particularly during the tenures of Presidents Biden and Trump.
+During his first term (2017–2021), Trump rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs.
+In his second term, the Trump administration has again taken a deregulatory approach.
+In contrast, the Biden administration adopted an aggressive pro-EV stance, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
+Table 3: Covariate effects (CE) from ordinal probit models. CE marked with $\times$ indicate parameters that are not statistically different from zero at 95% probability level.
+April 2021 May 2022 May–June 2023
+[…]
+EV Info 0.0851-0.0851 0.0388-0.0388 $0.0217$ $0.1022$ .... .... .... .... .... .... .... ....
+Env Better 0.1955-0.1955 0.0565-0.0565 $0.0970$ $0.1551$ .... .... .... .... .... .... .... ....
+EV Infra .... .... .... .... .... .... .... .... 0.1880-0.1880 0.0553-0.0553 $0.0526$ $0.1908$
+GACC (TM) $0.1971$ $0.0453$ 0.0945-0.0945 0.1478-0.1478 $0.3341$ $0.0133$ 0.1507-0.1507 0.1968-0.1968 $0.3170$ 0.0153-0.0153 0.1395-0.1395 0.1622-0.1622
+GACC (RA) $0.1180$ $0.0391$ 0.0490-0.0490 0.1082-0.1082 $0.1649$ $0.0267$ 0.0639-0.0639 0.1278-0.1278 $0.1282$ $0.0079$ 0.0486-0.0486 0.0875-0.0875
+EV Owner 0.1321-0.1321 0.0849-0.0849 $0.0051$ $0.2119$ 0.1749-0.1749 0.0794-0.0794 $0.0231$ $0.2312$ 0.1814-0.1814 0.0513-0.0513 $0.0421$ $0.1905$
+Age <50<50 0.0326-0.0326 0.0130-0.0130 $0.0099$ $0.0356$ 0.0410-0.0410 0.0090-0.0090 $0.0129$ $0.0371$ 0.0371-0.0371 0.0036-0.0036 $0.0126$ $0.0282$
+Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ 0.0082-0.0082 0.0240-0.0240 $0.0301$ $0.0027$ 0.0100-0.0100 0.0228-0.0228
+Grad & Above 0.0815-0.0815 0.0322-0.0322 $0.0276$ $0.0860$ 0.1139-0.1139 0.0240-0.0240 $0.0378$ $0.1001$ 0.1264-0.1264 0.0130-0.0130 $0.0445$ $0.0949$
+Some College 0.0373-0.0373 0.0152-0.0152 $0.0099$ $0.0426$ 0.0541-0.0541 0.0118-0.0118 $0.0156$ $0.0503$ 0.0517-0.0517 0.0052-0.0052 0.0166-0.0166 $0.0403$
+Married .... .... .... .... 0.0154-0.0154 0.0029-0.0029 $0.0050$ $0.0134$ $\times$ $\times$ $\times$ $\times$
+Metropolitan 0.0534-0.0534 0.0179-0.0179 $0.0191$ $0.0522$ 0.0662-0.0662 0.0099-0.0099 $0.0232$ $0.0528$ 0.0512-0.0512 0.0027-0.0027 $0.0183$ $0.0356$
+Midwest $0.0162$ $0.0061$ 0.0051-0.0051 0.0172-0.0172 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+South $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+West 0.0378-0.0378 0.0164-0.0164 $0.0101$ $0.0441$ 0.0457-0.0457 0.0109-0.0109 $0.0135$ $0.0431$ 0.0454-0.0454 0.0053-0.0053 $0.0148$ $0.0360$
+US Born $0.0416$ $0.0186$ 0.0106-0.0106 0.0497-0.0497 $0.0777$ $0.0217$ 0.0214-0.0214 0.0781-0.0781 $0.0551$ $0.0071$ 0.0178-0.0178 0.0444-0.0444
+Black $0.0614$ $0.0200$ 0.0220-0.0220 0.0594-0.0594 $0.0349$ $0.0058$ 0.0116-0.0116 0.0291-0.0291 $0.0460$ $0.0026$ 0.0160-0.0160 0.0326-0.0326
+Other Races $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Democrat 0.1211-0.1211 0.0578-0.0578 $0.0403$ $0.1386$ 0.1569-0.1569 0.0497-0.0497 $0.0513$ $0.1553$ 0.1882-0.1882 0.0353-0.0353 $0.0677$ $0.1558$
+Ind & Others 0.0808-0.0808 0.0310-0.0310 $0.0219$ $0.0898$ 0.0890-0.0890 0.0187-0.0187 $0.0266$ $0.0811$ 0.1091-0.1091 0.0093-0.0093 $0.0356$ $0.0827$
+[…]
+Income has a positive effect on EV purchase intent.
+Conditional on other factors, an additional \$50,000 in income changes the probability of strong interest (no interest) by 2.1–2.9 (2.2–3.6) percentage points.
 Our results find support in Zhao et al.
 (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals.
-The small covariate effect for income may reflect the growing availability of EVs across various market segments and the presence of subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
-Overall, the results show that interest in EVs varies based on environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographic factors.
-Some patterns likely reflect personal preferences, while others may reflect barriers such as limited awareness or charging access.
+The small covariate effect may reflect the growing availability of EVs across various market segments and the subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
+Overall, the results show that interest in EVs varies by environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographics.
+Some patterns likely reflect personal preferences; others may reflect barriers such as limited awareness or charging access.
 4.2 Results from Ordinal Quantile Models
-We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to the probabilities of outcomes at specific quantiles, conditional on the covariates.
-In particular, Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models along with the logarithm of marginal likelihood, estimated separately for each survey data.
-Table 5 reports the corresponding covariate effects.
-The covariate effects at the 50th percentile align closely with those from the ordinal probit model.
-However, across all surveys, the covariate effects at the 20th and 80th quantiles are noticeably different compared to those from the ordinal probit model.
-As such, we find compelling evidence that the impact of covariates on the intent to purchase EV is not uniform, but varies across the distribution of EV purchase intent.
-This highlights the value of quantile modeling for capturing these variations.
-Below, we emphasize the key differences across quantiles.
-While having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption.
-In the 2021 survey and at the 80th percentile of EV purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent to purchase compared to those who have read or heard little to nothing.
-At the 50th and 20th percentiles of EV purchase intent, the difference is 9.9 and 6.1 percentage points, respectively.
-The smaller covariate effect of prior exposure to EV information at the 20th percentile, suggests that simply raising general awareness may not be enough to persuade consumers with low purchase interest.
-In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may be effective in influencing purchase decisions.
-Similarly, the relationship between confidence that the US will build the necessary charging infrastructure (measured in the 2023 survey) and EV purchase intent varies across the distribution of EV purchase intent.
-At the 80th percentile, respondents who are confident in this development are 20.0 percentage points more likely to express strong intent to purchase compared to those who are not confident.
-The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile.
-Conversely, at the 80th percentile, those confident in infrastructure development (relative to those who are not) are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively.
-These findings suggest that confidence in infrastructure development strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent.
-Enhancing the affordability, accessibility, and reliability of charging infrastructure—and effectively communicating these efforts—can build consumer confidence and promote broader EV adoption.
-Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood, from April 2021, May 2022, and May–Jun 2023 survey data.
+We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates.
+To our knowledge, this aspect has not been examined in prior work on EVs.
+Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison (Maheshwari and Rahman, 2023).
+Note that the coefficients in Table 4 are not directly comparable to those in Table 2 for two main reasons: (a) the models operate on different scales due to differences in the variance of the error terms, and (b) they address distinct inferential goals-the ordinal probit model focuses on the average probability of outcomes, whereas the ordinal quantile model targets the probability of outcomes at specific quantiles.
+Table 5 reports the corresponding covariate effects for all variables whose parameters are statistically different from zero at 95% probability level.
+We see that covariate effects at the 50th percentile align closely with those from the ordinal probit model, but across all surveys, those at the 20th and 80th quantiles are noticeably different.
+This provides compelling evidence that the impact of covariates on purchase intent is not uniform but varies across the distribution of purchase intent, underscoring the value of quantile modeling in capturing such heterogeneity.
+Next, we emphasize the key differences across quantiles.
+Although having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption.
+In the 2021 survey and at the 80th percentile of purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent than those who have read or heard little to nothing.
+At the 50th and 20th percentiles of purchase intent, the difference is 9.9 and 6.1 percentage points, respectively.
+Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood ($\ln$ ML), from April 2021, May 2022, and May–Jun 2023 survey data.
 April 2021 May 2022 May–June 2023
@@
 Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
-Intercept $-3.60$ $0.25$ $-0.58$ $0.17$ $2.60$ $0.25$ $-2.15$ $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ $-2.81$ $0.30$ $-0.05$ $0.20$ $2.77$ $0.28$
-EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$.absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$.absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-EV Infra ..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-..absent...
-.
-$2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
+Intercept 3.60-3.60 $0.25$ 0.58-0.58 $0.17$ $2.60$ $0.25$ 2.15-2.15 $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ 2.81-2.81 $0.30$ 0.05-0.05 $0.20$ $2.77$ $0.28$
+EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$ .... .... .... .... .... .... .... .... .... .... .... ....
+Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$ .... .... .... .... .... .... .... .... .... .... .... ....
+EV Infra .... .... .... .... .... .... .... .... .... .... .... .... $2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
+GACC (TM) 2.81-2.81 $0.13$ 1.96-1.96 $0.09$ 2.49-2.49 $0.13$ 3.79-3.79 $0.16$ 2.63-2.63 $0.10$ 3.70-3.70 $0.14$ 4.00-4.00 $0.17$ 2.41-2.41 $0.10$ 3.45-3.45 $0.14$
+GACC (RA) 1.79-1.79 $0.10$ 1.24-1.24 $0.07$ 1.55-1.55 $0.11$ 1.94-1.94 $0.12$ 1.48-1.48 $0.08$ 2.03-2.03 $0.12$ 1.77-1.77 $0.12$ 1.15-1.15 $0.08$ 1.48-1.48 $0.12$
 EV Owner $2.80$ $0.13$ $1.76$ $0.10$ $2.91$ $0.16$ $2.94$ $0.16$ $1.83$ $0.11$ $3.31$ $0.20$ $2.65$ $0.16$ $1.99$ $0.11$ $3.08$ $0.19$
-GACC (TM) $-2.81$ $0.13$ $-1.96$ $0.09$ $-2.49$ $0.13$ $-3.79$ $0.16$ $-2.63$ $0.10$ $-3.70$ $0.14$ $-4.00$ $0.17$ $-2.41$ $0.10$ $-3.45$ $0.14$
-GACC (RA) $-1.79$ $0.10$ $-1.24$ $0.07$ $-1.55$ $0.11$ $-1.94$ $0.12$ $-1.48$ $0.08$ $-2.03$ $0.12$ $-1.77$ $0.12$ $-1.15$ $0.08$ $-1.48$ $0.12$
-Age$<50$ $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
-Female $-0.11$ $0.08$ $-0.02$ $0.05$ $-0.03$ $0.07$ $-0.36$ $0.10$ $-0.22$ $0.06$ $-0.33$ $0.09$ $-0.49$ $0.10$ $-0.31$ $0.07$ $-0.27$ $0.09$
-Grad & Above $1.42$ $0.14$ $-0.93$ $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
+Age<50<50 $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
+Female 0.11-0.11 $0.08$ 0.02-0.02 $0.05$ 0.03-0.03 $0.07$ 0.36-0.36 $0.10$ 0.22-0.22 $0.06$ 0.33-0.33 $0.09$ 0.49-0.49 $0.10$ 0.31-0.31 $0.07$ 0.27-0.27 $0.09$
+Grad & Above $1.42$ $0.14$ 0.93-0.93 $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
 Some College $0.66$ $0.14$ $0.47$ $0.09$ $0.57$ $0.12$ $1.08$ $0.16$ $0.53$ $0.10$ $0.49$ $0.13$ $0.92$ $0.16$ $0.42$ $0.10$ $0.53$ $0.13$
-Married ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
+Married .... .... .... .... .... .... $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
 Metropolitan $0.77$ $0.12$ $0.64$ $0.08$ $0.71$ $0.12$ $0.91$ $0.16$ $0.59$ $0.10$ $0.79$ $0.13$ $0.79$ $0.16$ $0.46$ $0.10$ $0.61$ $0.13$
-Midwest $-0.31$ $0.12$ $-0.19$ $0.08$ $-0.24$ $0.11$ $-0.09$ $0.15$ $-0.09$ $0.10$ $-0.11$ $0.14$ $-0.18$ $0.16$ $-0.15$ $0.10$ $-0.21$ $0.14$
-South $-0.21$ $0.10$ $-0.12$ $0.07$ $-0.17$ $0.10$ $0.00$ $0.13$ $-0.02$ $0.09$ $-0.06$ $0.13$ $-0.22$ $0.15$ $-0.22$ $0.09$ $-0.17$ $0.13$
+Midwest 0.31-0.31 $0.12$ 0.19-0.19 $0.08$ 0.24-0.24 $0.11$ 0.09-0.09 $0.15$ 0.09-0.09 $0.10$ 0.11-0.11 $0.14$ 0.18-0.18 $0.16$ 0.15-0.15 $0.10$ 0.21-0.21 $0.14$
+South 0.21-0.21 $0.10$ 0.12-0.12 $0.07$ 0.17-0.17 $0.10$ $0.00$ $0.13$ 0.02-0.02 $0.09$ 0.06-0.06 $0.13$ 0.22-0.22 $0.15$ 0.22-0.22 $0.09$ 0.17-0.17 $0.13$
 West $0.57$ $0.12$ $0.47$ $0.08$ $0.70$ $0.12$ $0.46$ $0.15$ $0.45$ $0.10$ $0.89$ $0.14$ $0.63$ $0.16$ $0.46$ $0.10$ $0.66$ $0.15$
-US Born $-0.63$ $0.12$ $-0.49$ $0.09$ $-0.80$ $0.14$ $-0.89$ $0.15$ $-0.67$ $0.10$ $-1.25$ $0.16$ $-0.91$ $0.16$ $-0.45$ $0.11$ $-0.67$ $0.15$
-Black $-1.23$ $0.15$ $-0.70$ $0.10$ $-0.56$ $0.14$ $-0.86$ $0.20$ $-0.25$ $0.14$ $-0.06$ $0.18$ $-0.94$ $0.16$ $-0.42$ $0.11$ $-0.36$ $0.16$
-Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ $-0.19$ $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ $-0.30$ $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
+US Born 0.63-0.63 $0.12$ 0.49-0.49 $0.09$ 0.80-0.80 $0.14$ 0.89-0.89 $0.15$ 0.67-0.67 $0.10$ 1.25-1.25 $0.16$ 0.91-0.91 $0.16$ 0.45-0.45 $0.11$ 0.67-0.67 $0.15$
+Black 1.23-1.23 $0.15$ 0.70-0.70 $0.10$ 0.56-0.56 $0.14$ 0.86-0.86 $0.20$ 0.25-0.25 $0.14$ 0.06-0.06 $0.18$ 0.94-0.94 $0.16$ 0.42-0.42 $0.11$ 0.36-0.36 $0.16$
+Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ 0.19-0.19 $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ 0.30-0.30 $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
 Democrat $2.00$ $0.11$ $1.33$ $0.08$ $1.94$ $0.12$ $2.30$ $0.14$ $1.45$ $0.09$ $1.90$ $0.13$ $2.49$ $0.14$ $1.71$ $0.10$ $2.40$ $0.14$
@@
 $\delta_{2}$ $1.53$ $0.02$ $0.95$ $0.02$ $1.35$ $0.02$ $1.46$ $0.02$ $0.84$ $0.02$ $1.16$ $0.02$ $1.52$ $0.02$ $0.93$ $0.02$ $1.19$ $0.02$
-$\ln$ ML $-12871$ $-12529$ $-12901$ $-8563$ $-8357$ $-8624$ $-8237$ $-7981$ $-8194$
-Table 5: Covariate effects (CE) from quantile analysis. CE for variables not statistically different from zero are marked with $\times$.
+$\ln$ ML 12871-12871 12529-12529 12901-12901 8563-8563 8357-8357 8624-8624 8237-8237 7981-7981 8194-8194
+Table 5: Covariate effects (CE) from quantile analysis. CE marked with $\times$ indicate parameters that are not statistically different from zero at 95% probability level.
 April 2021 Survey
 20th Quantile 50th Quantile 80th Quantile
-$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
-EV Info $-0.0722$ $-0.0421$ $0.0527$ $0.0615$ $-0.0792$ $-0.0374$ $0.0172$ $0.0995$ $-0.0722$ $-0.0453$ $-0.0089$ $0.1265$
-Env Better $-0.1958$ $-0.0191$ $0.1135$ $0.1014$ $-0.1771$ $-0.0730$ $0.1122$ $0.1379$ $-0.1384$ $-0.1127$ $0.0789$ $0.1722$
-EV Owner $-0.1357$ $-0.0896$ $0.0633$ $0.1620$ $-0.1212$ $-0.0757$ $-0.0113$ $0.2083$ $-0.1053$ $-0.0953$ $-0.0265$ $0.2270$
-GACC (TM) $0.2192$ $0.0097$ $-0.1217$ $-0.1072$ $0.1974$ $0.0635$ $-0.1203$ $-0.1406$ $0.1283$ $0.0779$ $-0.0575$ $-0.1487$
-GACC (RA) $0.1216$ $0.0312$ $-0.0780$ $-0.0747$ $0.1097$ $0.0534$ $-0.0613$ $-0.1018$ $0.0763$ $0.0421$ $-0.0124$ $-0.1059$
-Age $<50$ $-0.0237$ $-0.0106$ $0.0156$ $0.0187$ $-0.0303$ $-0.0134$ $0.0095$ $0.0343$ $-0.0265$ $-0.0164$ $0.0003$ $0.0427$
-Metropolitan $-0.0483$ $-0.0174$ $0.0318$ $0.0338$ $-0.0565$ $-0.0225$ $0.0237$ $0.0553$ $-0.0345$ $-0.0185$ $0.0030$ $0.0500$
-Midwest $0.0185$ $0.0080$ $-0.0123$ $-0.0141$ $0.0161$ $0.0068$ $-0.0053$ $-0.0177$ $0.0114$ $0.0066$ $-0.0003$ $-0.0177$
-South $0.0124$ $0.0056$ $-0.0083$ $-0.0098$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-West $-0.0327$ $-0.0168$ $0.0220$ $0.0275$ $-0.0382$ $-0.0176$ $0.0092$ $0.0467$ $-0.0311$ $-0.0200$ $-0.0015$ $0.0526$
-Black $0.0784$ $0.0256$ $-0.0519$ $-0.0520$ $0.0613$ $0.0244$ $-0.0259$ $-0.0598$ $0.0269$ $0.0141$ $-0.0014$ $-0.0396$
-Democrat $-0.1242$ $-0.0523$ $0.0808$ $0.0958$ $-0.1081$ $-0.0631$ $0.0431$ $0.1280$ $-0.0828$ $-0.0676$ $0.0070$ $0.1434$
+[…]
+EV Info 0.0722-0.0722 0.0421-0.0421 $0.0527$ $0.0615$ 0.0792-0.0792 0.0374-0.0374 $0.0172$ $0.0995$ 0.0722-0.0722 0.0453-0.0453 0.0089-0.0089 $0.1265$
+Env Better 0.1958-0.1958 0.0191-0.0191 $0.1135$ $0.1014$ 0.1771-0.1771 0.0730-0.0730 $0.1122$ $0.1379$ 0.1384-0.1384 0.1127-0.1127 $0.0789$ $0.1722$
+GACC (TM) $0.2192$ $0.0097$ 0.1217-0.1217 0.1072-0.1072 $0.1974$ $0.0635$ 0.1203-0.1203 0.1406-0.1406 $0.1283$ $0.0779$ 0.0575-0.0575 0.1487-0.1487
+GACC (RA) $0.1216$ $0.0312$ 0.0780-0.0780 0.0747-0.0747 $0.1097$ $0.0534$ 0.0613-0.0613 0.1018-0.1018 $0.0763$ $0.0421$ 0.0124-0.0124 0.1059-0.1059
+EV Owner 0.1357-0.1357 0.0896-0.0896 $0.0633$ $0.1620$ 0.1212-0.1212 0.0757-0.0757 0.0113-0.0113 $0.2083$ 0.1053-0.1053 0.0953-0.0953 0.0265-0.0265 $0.2270$
+Age <50<50 0.0237-0.0237 0.0106-0.0106 $0.0156$ $0.0187$ 0.0303-0.0303 0.0134-0.0134 $0.0095$ $0.0343$ 0.0265-0.0265 0.0164-0.0164 $0.0003$ $0.0427$
+Metropolitan 0.0483-0.0483 0.0174-0.0174 $0.0318$ $0.0338$ 0.0565-0.0565 0.0225-0.0225 $0.0237$ $0.0553$ 0.0345-0.0345 0.0185-0.0185 $0.0030$ $0.0500$
+Midwest $0.0185$ $0.0080$ 0.0123-0.0123 0.0141-0.0141 $0.0161$ $0.0068$ 0.0053-0.0053 0.0177-0.0177 $0.0114$ $0.0066$ 0.0003-0.0003 0.0177-0.0177
+South $0.0124$ $0.0056$ 0.0083-0.0083 0.0098-0.0098 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+West 0.0327-0.0327 0.0168-0.0168 $0.0220$ $0.0275$ 0.0382-0.0382 0.0176-0.0176 $0.0092$ $0.0467$ 0.0311-0.0311 0.0200-0.0200 0.0015-0.0015 $0.0526$
+Black $0.0784$ $0.0256$ 0.0519-0.0519 0.0520-0.0520 $0.0613$ $0.0244$ 0.0259-0.0259 0.0598-0.0598 $0.0269$ $0.0141$ 0.0014-0.0014 0.0396-0.0396
+Democrat 0.1242-0.1242 0.0523-0.0523 $0.0808$ $0.0958$ 0.1081-0.1081 0.0631-0.0631 $0.0431$ $0.1280$ 0.0828-0.0828 0.0676-0.0676 $0.0070$ $0.1434$
 May 2022 Survey
-EV Owner $-0.1782$ $-0.0723$ $0.0880$ $0.1624$ $-0.1573$ $-0.0695$ $0.0121$ $0.2147$ $-0.1465$ $-0.1110$ $-0.0071$ $0.2646$
-GACC (TM) $0.3425$ $-0.0568$ $-0.1515$ $-0.1341$ $0.3304$ $0.0226$ $-0.1778$ $-0.1752$ $0.2471$ $0.0963$ $-0.1459$ $-0.1975$
-GACC (RA) $0.1541$ $0.0026$ $-0.0790$ $-0.0777$ $0.1527$ $0.0497$ $-0.0852$ $-0.1171$ $0.1269$ $0.0392$ $-0.0308$ $-0.1352$
-Age $<50$ $-0.0362$ $-0.0066$ $0.0199$ $0.0229$ $-0.0390$ $-0.0085$ $0.0130$ $0.0344$ $-0.0358$ $-0.0167$ $0.0072$ $0.0453$
-Metropolitan $-0.0684$ $-0.0054$ $0.0360$ $0.0378$ $-0.0643$ $-0.0109$ $0.0258$ $0.0493$ $-0.0483$ $-0.0165$ $0.0114$ $0.0535$
-Black $0.0642$ $0.0060$ $-0.0342$ $-0.0359$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Democrat $-0.1719$ $-0.0295$ $0.0927$ $0.1087$ $-0.1437$ $-0.0574$ $0.0601$ $0.1410$ $-0.0996$ $-0.0659$ $0.0217$ $0.1438$
+GACC (TM) $0.3425$ 0.0568-0.0568 0.1515-0.1515 0.1341-0.1341 $0.3304$ $0.0226$ 0.1778-0.1778 0.1752-0.1752 $0.2471$ $0.0963$ 0.1459-0.1459 0.1975-0.1975
+GACC (RA) $0.1541$ $0.0026$ 0.0790-0.0790 0.0777-0.0777 $0.1527$ $0.0497$ 0.0852-0.0852 0.1171-0.1171 $0.1269$ $0.0392$ 0.0308-0.0308 0.1352-0.1352
+EV Owner 0.1782-0.1782 0.0723-0.0723 $0.0880$ $0.1624$ 0.1573-0.1573 0.0695-0.0695 $0.0121$ $0.2147$ 0.1465-0.1465 0.1110-0.1110 0.0071-0.0071 $0.2646$
+Age <50<50 0.0362-0.0362 0.0066-0.0066 $0.0199$ $0.0229$ 0.0390-0.0390 0.0085-0.0085 $0.0130$ $0.0344$ 0.0358-0.0358 0.0167-0.0167 $0.0072$ $0.0453$
+Metropolitan 0.0684-0.0684 0.0054-0.0054 $0.0360$ $0.0378$ 0.0643-0.0643 0.0109-0.0109 $0.0258$ $0.0493$ 0.0483-0.0483 0.0165-0.0165 $0.0114$ $0.0535$
+Black $0.0642$ $0.0060$ 0.0342-0.0342 0.0359-0.0359 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Democrat 0.1719-0.1719 0.0295-0.0295 $0.0927$ $0.1087$ 0.1437-0.1437 0.0574-0.0574 $0.0601$ $0.1410$ 0.0996-0.0996 0.0659-0.0659 $0.0217$ $0.1438$
 May–June 2023 Survey
-EV Infra $-0.1863$ $-0.0428$ $0.0956$ $0.1335$ $-0.1827$ $-0.0603$ $0.0616$ $0.1813$ $-0.1452$ $-0.0730$ $0.0187$ $0.1995$
-EV Owner $-0.1684$ $-0.0350$ $0.0727$ $0.1307$ $-0.1842$ $-0.0542$ $0.0459$ $0.1925$ $-0.1580$ $-0.0817$ $0.0183$ $0.2215$
-GACC (TM) $0.3323$ $-0.0635$ $-0.1405$ $-0.1283$ $0.3103$ $-0.0250$ $-0.1426$ $-0.1427$ $0.2512$ $0.0486$ $-0.1423$ $-0.1575$
-GACC (RA) $0.1318$ $-0.0065$ $-0.0592$ $-0.0661$ $0.1238$ $0.0132$ $-0.0570$ $-0.0799$ $0.0975$ $0.0183$ $-0.0329$ $-0.0830$
-Age $<50$ $-0.0296$ $-0.0007$ $0.0131$ $0.0173$ $-0.0363$ $-0.0034$ $0.0139$ $0.0258$ $-0.0349$ $-0.0082$ $0.0102$ $0.0329$
-Metropolitan $-0.0568$ $0.0013$ $0.0250$ $0.0304$ $-0.0511$ $-0.0026$ $0.0204$ $0.0333$ $-0.0408$ $-0.0067$ $0.0129$ $0.0345$
-Black $0.0665$ $-0.0010$ $-0.0296$ $-0.0359$ $0.0462$ $0.0029$ $-0.0182$ $-0.0309$ $0.0239$ $0.0039$ $-0.0069$ $-0.0209$
-Democrat $-0.1832$ $-0.0084$ $0.0851$ $0.1066$ $-0.1815$ $-0.0384$ $0.0798$ $0.1401$ $-0.1393$ $-0.0660$ $0.0491$ $0.1562$
-Beliefs about EVs’ environmental benefits (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles.
-Among respondents at the 80th percentile, those who believe EVs are better for the environment are 17.2 percentage points more likely to express strong purchase intent compared to those at the same percentile who do not share this belief.
+EV Infra 0.1863-0.1863 0.0428-0.0428 $0.0956$ $0.1335$ 0.1827-0.1827 0.0603-0.0603 $0.0616$ $0.1813$ 0.1452-0.1452 0.0730-0.0730 $0.0187$ $0.1995$
+GACC (TM) $0.3323$ 0.0635-0.0635 0.1405-0.1405 0.1283-0.1283 $0.3103$ 0.0250-0.0250 0.1426-0.1426 0.1427-0.1427 $0.2512$ $0.0486$ 0.1423-0.1423 0.1575-0.1575
+GACC (RA) $0.1318$ 0.0065-0.0065 0.0592-0.0592 0.0661-0.0661 $0.1238$ $0.0132$ 0.0570-0.0570 0.0799-0.0799 $0.0975$ $0.0183$ 0.0329-0.0329 0.0830-0.0830
+EV Owner 0.1684-0.1684 0.0350-0.0350 $0.0727$ $0.1307$ 0.1842-0.1842 0.0542-0.0542 $0.0459$ $0.1925$ 0.1580-0.1580 0.0817-0.0817 $0.0183$ $0.2215$
+Age <50<50 0.0296-0.0296 0.0007-0.0007 $0.0131$ $0.0173$ 0.0363-0.0363 0.0034-0.0034 $0.0139$ $0.0258$ 0.0349-0.0349 0.0082-0.0082 $0.0102$ $0.0329$
+Metropolitan 0.0568-0.0568 $0.0013$ $0.0250$ $0.0304$ 0.0511-0.0511 0.0026-0.0026 $0.0204$ $0.0333$ 0.0408-0.0408 0.0067-0.0067 $0.0129$ $0.0345$
+Black $0.0665$ 0.0010-0.0010 0.0296-0.0296 0.0359-0.0359 $0.0462$ $0.0029$ 0.0182-0.0182 0.0309-0.0309 $0.0239$ $0.0039$ 0.0069-0.0069 0.0209-0.0209
+Democrat 0.1832-0.1832 0.0084-0.0084 $0.0851$ $0.1066$ 0.1815-0.1815 0.0384-0.0384 $0.0798$ $0.1401$ 0.1393-0.1393 0.0660-0.0660 $0.0491$ $0.1562$
+Similarly, the relationship between confidence about building charging infrastructure (measured in the 2023 survey) and purchase intent varies across the distribution of the latter.
+At the 80th percentile, confident respondents are 20.0 percentage points more likely to express strong purchase intent than those who are not confident.
+The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile.
+Conversely, at the 80th percentile, those who are confident are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively.
+Beliefs about EVs’ environmental benefits relative to gas-powered vehicles (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles.
+Among respondents at the 80th percentile, those who believe EVs are environmentally better than gas-powered vehicles are 17.2 percentage points more likely to express strong purchase intent than those at the same percentile who do not share this belief.
 The corresponding differences are 13.8 percentage points at the median and 10.1 percentage points at the 20th percentile.
-Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 percentage points less likely at the median and 19.6 percentage points less likely at the 20th percentile.
-This finding suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters while easing skepticism among consumers with lower purchase interest; this indicates that highlighting benefits such as local air quality improvements could broaden support for EV adoption.
+Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 and 19.6 percentage points less likely at the median and 20th percentile, respectively.
+Regarding views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong purchase interest, with this becoming more pronounced at higher percentiles.
+In the 2021 survey, those at the 80th percentile of purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent than those at the same percentile who do not.
+The gap is 14.1 and 10.7 percentage points at the median and 20th percentile, respectively.
+Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 and 21.9 percentage points at the median and 20th percentile, respectively.
+More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much approximately 33 and 25 percentage points more likely to express no interest in EVs at the 20th percentile and 80th percentile, respectively.
 The influence of prior EV ownership also varies across quantiles.
 At the 80th percentile, respondents with prior EV ownership are 22.7 percentage points more likely to express strong purchase intent compared to those without prior ownership.
-This difference decreases to 20.8 percentage points at the median and 16.2 percentage points at the 20th percentile.
+This difference decreases to 20.8 and 16.2 percentage points at the median and 20th percentile, respectively.
 A similar pattern is observed in both the 2022 and 2023 surveys.
-In the 2022 survey, the probability of strong interest increases by 26.5, 21.5, and 16.2 percentage points at the 80th, 50th, and 20th quantiles, respectively.
-For the same quantiles, the covariate effects for strong interest in the 2023 survey are 22.1, 19.2, and 13.1 percentage points.
-When it comes to views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong interest in purchasing an EV, with this negative effect becoming more pronounced at higher percentiles.
-In the 2021 survey, those at the 80th percentile of EV purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent compared to those at the same percentile who do not.
-The gap is 14.1 percentage points at the median and 10.7 percentage points at the 20th percentile.
-Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 percentage points at the median and 21.9 percentage points at the 20th percentile.
-More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much on climate issues being approximately 33 percentage points more likely to express no interest in EVs at the 20th percentile and 25 percentage points more likely at the 80th percentile.
-These patterns suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonates across ideological lines.
-Rather than positioning EVs primarily as a climate solution, emphasizing other benefits—such as long-term cost savings and air quality improvements—may be more effective in engaging a broader range of potential buyers.
-Moreover, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns, particularly for consumers who are wary of climate-focused policies.
-Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments.
-Legend: ΔPr(Not at all $\Delta\Pr(\textrm{Not at all likely})$:  ; ΔPr(Not too $\Delta\Pr(\textrm{Not too likely})$:   ; ΔPr(Somewhat $\Delta\Pr(\textrm{Somewhat likely})$:   ; ΔPr(Very $\Delta\Pr(\textrm{Very likely})$:  .
+Across surveys and quantiles, prior EV ownership is the strongest predictor for EV purchase intent.
 Geographic and demographic factors show smaller but consistent relationships with EV purchase intent across all surveys.
@@
 Finally, the relationship between political party affiliation and EV purchase intent varies considerably across quantiles.
-Results from the 2021 survey show that at the 80th percentile Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points among respondents at the 20th percentile.
-Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, while the difference was 12.4 percentage points at the 20th percentile.
+Results from the 2021 survey show that at the 80th percentile, Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points at the 20th percentile.
+Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, but the difference is 12.4 percentage points at the 20th percentile.
 These patterns remain consistent in more recent surveys.
-The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, while at lower quantiles, they are far less likely to reject EVs outright.
-Emphasizing benefits that resonate across political ideologies—such as long-term cost savings and energy independence—may help broaden EV appeal across the political spectrum.
-To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents the covariate effects for some selected variables from a series of ordinal quantile regression models, covering the 5th to the 95th percentiles of EV purchase intent in 5-percentile increments.
+The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, and at lower quantiles, they are far less likely to reject EVs outright.
+Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments.
+Legend: $\Delta\Pr(\textrm{Not at all likely})$:  ; $\Delta\Pr(\textrm{Not too likely})$:   ; $\Delta\Pr(\textrm{Somewhat likely})$:   ; $\Delta\Pr(\textrm{Very likely})$:  .
+To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents these for some selected variables from a series of ordinal quantile regression models, covering the 5th to 95th percentiles of EV purchase intent in 5-percentile increments.
 Each panel corresponds to a different covariate; with the solid gray, dashed gray, dashed black, and solid black lines representing the change in probability of reporting “not likely at all,” “not too likely,” “somewhat likely,” and “very likely,” respectively, for a given change in covariate across the quantiles.
-The figure clearly illustrates the substantial differences in covariate effects across the distribution of EV purchase intent.
-For example, the solid black line in these panels indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have previously owned an EV, identify as Democrats, or are confident in the United States’ ability to develop the necessary infrastructure are much more likely to report strong interest in purchasing EVs compared to those at lower quantiles.
-Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest in buying EVs at lower quantiles and markedly less likely to report strong interest in buying EVs at higher quantiles.
-Overall, these results highlight the need for tailored strategies to address the barriers to EV adoption for different population segments.
+The figure clearly illustrates the substantial differences in covariate effects across the distribution.
+For example, the solid black line indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have owned one, identify as Democrats, or are confident in the US ability to develop infrastructure are much more likely to report strong purchase interest compared to those at lower quantiles.
+Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest at lower quantiles and markedly less likely to report strong interest at higher quantiles.
+5 Discussion
+Our findings demonstrate how targeted strategies may be able to accelerate EV adoption.
+Exposure to EV information is associated with a 10.2 percentage point higher probability of reporting strong intent, with larger differences among respondents at the upper end of the intent distribution.
+It is also associated with an 8.5 point lower probability of reporting no intent.
+However, only 37.7% of respondents report high exposure, pointing to an opportunity for broader outreach through media, industry, and education.
+The smaller covariate effect of prior exposure at the 20th percentile suggests that simply raising awareness may not persuade consumers with low purchase interest.
+In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may influence purchase decisions.
+Belief that EVs are better for the environment than gas-powered vehicles is associated with a 15.5 percentage point higher probability of reporting strong intent, with larger differences at the upper end of the distribution.
+It is also associated with a 19.5 point lower probability of reporting no interest, with larger differences at the lower end of the distribution.
+This pattern suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters and reduces skepticism among consumers with lower purchase interest.
+In the 2021 survey, 71.8% of respondents expressed this belief.
+Emphasizing environmental benefits—such as improved air quality—may resonate with both likely adopters and more skeptical respondents, broadening support for EV adoption.
+Confidence in charging infrastructure is one of the strongest predictors of purchase intent.
+Respondents who express confidence that the US will build sufficient infrastructure are 19.1 percentage points more likely to report strong intent and 18.8 points less likely to report no interest.
+Quantile estimates suggest that confidence both strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent.
+Despite the importance of this factor, only 17.5% of respondents in the 2023 survey reported high confidence.
+Policymakers and industry can address this gap through continued investment in charging infrastructure-such as the National Electric Vehicle Infrastructure (NEVI) program-to improve affordability, accessibility, and reliability, as well as through effective communication of these efforts and of a long-term vision for the charging network.
+Belief that the government is doing too much to address climate change is associated with a 14.8–19.7 point lower probability of strong intent and a 19.7–33.4 point higher probability of no interest.
+In the 2023 survey, 24.1% of respondents held this view.
+These findings suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonate across ideological lines.
+Rather than positioning EVs primarily as a climate solution, highlighting other benefits—such as long-term cost savings and air quality improvements—may engage a broader range of potential buyers.
+Our finding that Democrats are more likely to express stronger interest in EVs, further highlights the role of political identity and the value of emphasizing non-partisan benefits to broaden appeal.
+Moreover, for consumers wary of climate-focused policies, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns.
+The results for other covariates point to further implications.
+Prior EV ownership is the strongest predictor of purchase intent, underscoring the importance of reducing barriers for first-time buyers.
+Potential options include subsidies, test-drive programs, and community-based education.
+Younger, urban, educated, and higher-income individuals report greater interest in EVs, suggesting that targeting information or advertising to such individuals may help raise adoption rates.
+By contrast, lower interest among women, US-born, and Black respondents may reflect cultural, structural, or informational barriers that could be addressed through inclusive policy design and outreach.
 6 Conclusion
-EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts.
-However, government support has varied between administrations.
-The Obama administration facilitated EV adoption by introducing federal tax credits for EV purchases and strengthening fuel economy and emissions standards, which incentivized automakers to develop electric models.
-In contrast, the Trump administration’s first term prioritized deregulation and rolled back these standards.
-The Biden administration implemented stricter emissions regulations and allocated substantial funding for EV infrastructure through the Bipartisan Infrastructure Law.
-A second Trump administration has signaled a likely return to deregulation, though specific policies remain unclear.
-While policies and regulatory standards continue to evolve with changing administrations, the future of electric vehicles depends largely on the interest of consumers, which plays a crucial role in shaping national policies.
-Public opinion has a well-documented influence on US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has systematically analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of the covariates.
+EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts (Bhat et al., 2025).
+However, federal support has varied along partisan lines, with Democratic administration promoting EV adoption and Republican administration prioritizing deregulation and rolling back environmental standards.
+Amid these policy shifts, consumer interest remains a key driver of both market trends and national policy.
+Although public opinion is known to influence US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of covariates.
 In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
-This study addresses these gaps by examining public attitudes toward EVs using three nationally representative surveys conducted by the Pew Research Center from 2021-2023.
-Given the ordinal nature of the response data, we employ both ordinal probit regression and ordinal quantile regression to model EV purchase intent as a function of four novel variables not considered in existing studies, prior EV ownership, political affiliation, and demographic characteristics such as age, gender, income, and geographic location.
-We adopt the Bayesian approach to modeling and estimation, as it offers the entire posterior distribution of the parameters, enables direct probability statements, and allows us to account for uncertainty in observations and parameters when evaluating the effects of covariates.
-The results from ordinal probit models indicate that respondents who say they have read or heard a lot about EVs, perceive them as environmentally beneficial, have previously owned an EV or hybrid vehicle, and are confident in the US developing charging infrastructure are more likely to express strong interest in purchasing an EV.
-Other factors positively influencing the intention to purchase an EV include being under 50 years old, having a higher level of education, living in metropolitan areas, residing in the West (compared to the Northeast), identifying with the Democratic Party (rather than the Republican Party), and having a higher income.
-In contrast, those who believe government climate efforts are excessive or adequate, are female, live in the Midwest, were born in the US, or are Black (as opposed to White) are less likely to express strong interest and more likely to report no interest in EVs.
-Results from covariate effects reveal that prior EV ownership, confidence in EV infrastructure, belief in EVs’ environmental benefits, attitudes toward government climate policy, and political affiliation have the strongest influence on individuals’ decision to purchase EVs.
-In addition, results from ordinal quantile analysis show that the relationship between these factors and EV purchase intent varies across the conditional distribution of the intention to purchase EVs.
-Respondents who are knowledgeable about EVs, believe in their environmental benefits, are confident in US charging infrastructure, have prior EV ownership, or identify as Democrats are especially likely to express strong purchase intent at higher quantiles of EV purchase intent.
-Similarly, those who believe EVs are environmentally beneficial, have confidence in infrastructure development, or identify as Democrats are particularly less likely to express no interest at lower quantiles.
-In contrast, respondents who believe the government is doing too much to address climate change are particularly more likely to express no interest at lower quantiles and particularly less likely to express strong interest at higher quantiles.
-Overall, the quantile models highlight significant variation in covariate effects across the distribution of EV purchase intent; thereby providing a more nuanced understanding than the ordinal probit model.
-The findings demonstrate how targeted strategies may be able to help accelerate EV adoption.
-For consumers with low purchase intent (i.e., at lower quantiles), general awareness campaigns may have limited impact; instead, emphasizing practical aspects such as long-term cost savings, availability and reliability of charging infrastructure, and financial incentives could be more effective.
-For more interested consumers (i.e., at higher quantiles), reinforcing beliefs about environmental benefits can transform a likely customer to an actual EV owner.
-Confidence in charging infrastructure plays a dual role, both strengthening commitment among likely adopters and reducing outright rejection among skeptics, highlighting the potential value of visible investments and clear public communication about infrastructure progress.
-The results also indicate that political identity and views on climate policy are closely related to EV purchase intent, suggesting that framing EVs in terms of widely appealing features such as financial savings and energy independence—rather than solely as a climate solution—could resonate with a broader range of potential buyers.
-Moreover, for consumers wary of government climate policy, information provided by automakers, dealerships, and industry groups may carry more credibility than government-led campaigns.
-Overall, the empirical findings of our study provide useful insights for policymakers, industry stakeholders, and advertisers seeking to bolster adoption.
+This study addresses these gaps by analyzing public attitudes toward EVs using three nationally representative Pew Research Center surveys from 2021-2023.
+We take the Bayesian route and model EV purchase intent with ordinal probit and ordinal quantile regression, incorporating novel predictors, prior EV ownership, political affiliation, and demographic factors.
+Results from the ordinal probit models show that strong interest in purchasing an EV is associated with greater exposure to EV information, belief in their environmental benefits, prior EV or hybrid ownership, and confidence in charging infrastructure.
+Additional positive predictors include being under 50, higher education and income, metropolitan residence, living in the West, and Democratic Party affiliation.
+Conversely, lower interest is linked to skepticism about government climate efforts, identifying as female or Black, living in the Midwest, and US-born status.
+Ordinal quantile analysis reveals that the effects of key factors on EV purchase intent vary across the distribution.
+Respondents who are knowledgeable about EVs, view them as environmentally beneficial, have confidence in charging infrastructure, own an EV, or identify as Democrats are more likely to express strong interest at higher quantiles of EV purchase intent.
+These same factors are associated with lower likelihood of no interest at lower quantiles.
+Conversely, those who believe the government is overreaching on climate policy are more likely to express no interest at lower quantiles and less likely to show strong interest at higher ones.
+These findings highlight how covariate effects differ across levels of purchase intent, thereby providing a more nuanced understanding than the ordinal probit model.
+The findings offer valuable guidance for policymakers, industry stakeholders, and advertisers seeking to bolster EV adoption.
+Confidence in charging infrastructure is one of the strongest predictors of purchase intent—associated with greater interest among likely adopters and reduced rejection among skeptics—yet less than 20% respondents report high confidence.
+Infrastructure investment, along with clearer communication about availability, reliability and expansion plans, may help close this gap.
+Exposure to EV information is also linked to stronger intent and reduced disinterest, though only 37.7% of respondents report high exposure—highlighting the value of broader outreach through media, industry, and education.
+Belief that EVs are environmentally superior to gas-powered vehicles is similarly associated with higher intent and lower disinterest, with effects varying across the intent distribution; messaging focused on benefits such as improved air quality may therefore be effective.
+Perceptions that the government is doing too much on climate change are negatively associated with intent, suggesting that outreach emphasizing cost, convenience, or energy security—particularly from non-government sources—may have broader appeal.

### v3

@@
 Decarbonization; Electric vehicle; Ordinal probit; Pew Research Center; Quantile regression; Technology adoption.
-journal: Transportation Research Part A
+journal: Transportation Research Part A: Policy and Practice
 1 Introduction
-Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024).
-Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets (Bhat et al., 2025).
-To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives (Jenn et al., 2018; Mekky and Collins, 2024).
+Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half [18].
+Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets [6].
+Over the past decade, governments worldwide, including the US, have implemented a range of policies and incentives aimed at encouraging EV adoption [28, 43].
 Key US measures have included nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program.
-As of late 2025, several of these EV-focused policies have been rolled back under the Trump administration..
-Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024).
+As of early 2026, several of these EV-focused policies have been rolled back under the Trump administration..
+Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) [23].
 The trajectory of US adoption is also increasingly uncertain.
-In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024).
+In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand [15].
 Successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations.
 The Biden administration promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its broader climate agenda.
-In contrast, the Trump administration has pursued a deregulatory shift: as of late 2025, the federal EV tax credits are being phased out and the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked.
+In contrast, the Trump administration has pursued a deregulatory shift: as of early 2026, federal EV tax credits have been eliminated, the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked.
 [^2] For details on the EV tax credit phase-out, see Section 70502 of the One Big Beautiful Bill Act.
@@
 These changing political priorities reflect tensions between environmental regulations, industry interests, and fiscal considerations in promoting EV adoption.
-Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024).
+Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges [4].
 The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and limit consumer choice.
-Ultimately, the adoption of EVs depends on consumer interest, which is likely to be shaped by multiple factors, such as cost, convenience, perception of environmental benefits, and awareness of these trade-offs (Rezvani et al., 2015).
-Although EVs often have higher upfront costs due to expensive batteries, they can offer long-term savings as electric motors require less maintenance (Liu et al., 2021).
-However, the overall cost of ownership can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024).
-Access to reliable charging infrastructure also remains an important consideration (Chakraborty et al., 2019, 2022).
-Limited availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel.
-Environmental concerns also affect consumer attitudes, but perceptions may vary, given the regional differences in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
-Recognizing the importance of public interest in shaping the future of this market, we examine how stated purchase intent in the US relates to the following four key factors: exposure to EV information, perceptions of their environmental benefits (compared to gas-powered vehicles), perceptions of government action to mitigate climate change, and confidence in infrastructure development.
-At the same time, we control for prior ownership, and commonly used demographic variables such as age, gender, education, political affiliation, and geographic location.
-The data for these four focal factors, along with the dependent and control variables, are drawn from three nationally representative opinion surveys conducted by the Pew Research Center over multiple years.
-Our empirical analysis consists of two stages.
-First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how these factors relate to an ordinal measure of EV purchase intent.
-We present the covariate effects on the average probabilities of reporting each level of interest, an aspect rarely explored in the EV literature.
-Secondly, and perhaps more importantly, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how these factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined.
-The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
-Our paper makes three key contributions.
-First, we examine four unique factors (while controlling for a host of demographic variables) to understand how informational, perceptual, and trust-related factors influence the likelihood and intensity of stated EV purchase intent in the US, and how do these effects vary across different levels of consumer interest.
-Prior studies of EV purchase intent have considered related factors, but these have either been distinct from the ones we study or investigated only in small samples.
-The factors we examine have not been analyzed jointly, on large samples, or in the US context.
-The first factor is confidence in future charging infrastructure development.
-While earlier work has emphasized the role of existing infrastructure (Carley et al., 2013; Tiwari et al., 2020), purchase intent is also likely to depend on expectations about future network expansion, particularly when the current system remains incomplete.
-The second factor is perceptions of government climate efforts.
-Past studies have examined links between purchase intent and political affiliation (e.g., Mamkhezri (2025); Sintov et al.
-(2020)), but not views on government climate action.
-Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase—especially if they are less aware of other benefits.
+Amid fluctuating incentives and persistent regulatory uncertainty, EV adoption ultimately depends on consumer interest.
+Accordingly, the study of public opinion on EV adoption is particularly important, as public opinion is known to influence US public policy even in the presence of competing interests and partisan conflict [9, 50].
+While the literature on EV adoption is extensive, important knowledge gaps remain.
+We organize these gaps, and our contributions to addressing them, along three dimensions: (a) the covariates (independent variables or regressors) considered, (b) the scale and representativeness of the data, and (c) the statistical models used to examine the relationship between EV adoption and these covariates.
+Within the first dimension, EV adoption is typically modeled as a function of demographic characteristics–such as age, gender, education, race, political affiliation, and income–together with factors related to cost [39, 56], convenience and charging infrastructure [12, 11], perception of environmental benefits [20, 35], and awareness of these trade-offs [47].
+In this context, any determinant that plausibly affects EV adoption but has not been systematically examined represents an opportunity to deepen our understanding of adoption behavior.
+Building on this literature, we examine EV adoption using standard demographic controls and introduce four additional factors: confidence in building charging stations and infrastructure, perception of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits relative to gasoline powered vehicles.
+The first factor–confidence in development of future charging infrastructure–is crucial for EV adoption because purchase intent is likely to depend on expectations about future network expansion, particularly when the current system remains incomplete.
+This stands in contrast to existing studies such as Carley et al.
+[10] and Tiwari et al.
+[55], which emphasize the role of existing infrastructure.
+The second factor concerns perceptions of government climate efforts.
+While previous research has documented links between EV purchase intent and political affiliation (e.g., Mamkhezri [42], Sintov et al.
+[53]), it has not paid attention to individuals’ assessments of government climate action itself.
+Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase an EV—particularly if they are less aware of EVs’ ancillary benefits.
 The third factor is exposure to EV-related information.
-Previous research has examined the effects of information on total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), and peer effects (Zhao et al., 2022).
-We build on this by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference.
+Previous research has examined the effects of information on total cost of ownership [17], media coverage [49], and peer effects [58].
+We extend these works by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference.
 The fourth factor is perceptions of EVs’ environmental benefits.
-Several studies have considered consumers’ pro-environmental attitudes (Hu et al., 2023; Carley et al., 2013; Krause et al., 2013; Khatua et al., 2023; Scherrer, 2023; Mamkhezri, 2025) and concern about reputation (Buhmann and Criado, 2023).
+Several studies have considered consumers’ pro-environmental attitudes [22, 10, 33, 31, 49, 42] and reputational concerns [8].
 However, research on perceptions of EV’s environmental benefits relative to gasoline powered vehicles is limited.
 Lashari et al.
-(2021) and Jia et al.
-(2025) examine this factor in a small sample, whereas our analysis situates it in a broader US context.
+[36] and Jia et al.
+[29] examine this factor in a small sample, whereas our analysis situates it in a broader US context.
 [^3] Studies have also examined other factors such as range anxiety, performance concerns, and consumer incentives (See Carley et al.
-(2013); Tiwari et al.
-(2020); Zhao et al.
-(2022); Stekelberg and Vance (2024); Mamkhezri (2025) among others).
-A separate strand of research examines determinants of actual EV sales, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), peer effects (Chakraborty et al., 2022), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023).
-Our findings reveal sizable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
-Our second contribution is the use of large, nationally representative datasets, which stands in contrast to the predominantly small-scale, non-representative surveys commonly used in prior research on EV purchase intent.
+[10], Tiwari et al.
+[55], Zhao et al.
+[58], Stekelberg and Vance [54], Mamkhezri [42] among others).
+A separate strand of research examines determinants of actual EV sales, including tax credits and rebates [16, 28, 14, 51], charging infrastructure [37], peer effects [11], and automakers’ responses to environmental policies [1, 19, 40].
+Turning to the second dimension–the scale and representativeness of data–we observe that much of the existing research on EV purchase intent relies on single-year surveys, non-representative samples, or relatively small datasets.
 For example, Tiwari et al.
-(2020), Hu et al.
-(2023), and Mamkhezri (2025) analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the United States, respectively.
+[55], Hu et al.
+[22], and Mamkhezri [42] analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the US, respectively.
 [^4] Although some recent studies (e.g., Pani et al.
-(2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it.
+[44] and Ruan and Lv [48]) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it.
 Singh et al.
-(2020) provide a meta-analysis of 211 peer-reviewed research articles 2009–2019, identifying key factors influencing public opinion of EVs.
-Notably, most such studies are based on single-year, non-representative, or small samples.
-We overcome these limitations by using three nationally representative US opinion polls, with significantly larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023.
-This allows for greater statistical power and more robust generalizability of our findings.
-Our third-and perhaps most significant-contribution lies in applying advanced modeling and estimation techniques to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent.
-Research has used various methods, such as linear regression to assess the effects of personal beliefs, preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation (Buhmann and Criado, 2023); mixed logit models to explore attitudes toward various EV policy incentives, employment changes, and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers, such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022).
-A key limitation of these approaches is that they cannot capture heterogeneity in covariate effects across the full distribution of intent.
-We address this by applying ordinal quantile analysis to EV purchase intent, an approach not previously used in this context.
-In contrast to prior work that focuses on average effects, this approach captures heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase).
-This yields more detailed evidence on how different factors shape adoption intent.
+[52] provide a meta-analysis of 211 peer-reviewed research articles between 2009–2019, identifying key factors influencing public opinion of EVs.
+We overcome the limitations associated with small data by using three nationally representative US opinion polls, with substantially larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023.
+This allows for greater statistical power and more robust generalizability of our findings
+Finally, along the third dimension–use of statistical models to analyze EV adoption–we note that most studies have relied on linear regression to assess the effects of personal beliefs, preferences [10, 33] and positive media coverage [49].
+Some studies have employed binary probit and logit models to examine motivational factors such as reputation [8]; while others have used mixed logit models to analyze attitudes toward various EV policy incentives, employment changes, and electricity costs [42].
+However, in much of this literature, covariate effects surprisingly remains unreported, limiting the interpretability of the estimated relationships.
+Beyond these approaches, structural equation models have also been used to examine adoption barriers [55], peer effects [58], and policy incentives [22]; while topic modeling and sentiment analysis have been applied to track temporal shifts in online sentiment toward EVs [48].
+Nonetheless, all the above mentioned modeling frameworks cannot capture the heterogeneity in covariate effects across the full distribution of intent, which is necessary for a careful and deeper understanding of EV adoption.
+We employ a Bayesian approach and contribute to the modeling of EV adoption literature in two key ways.
+First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how our four factors and demographic characteristics relate to an ordinal measure of EV purchase intent.
+We report covariate effects on the average probabilities of each response category, an aspect that has received limited attention in prior studies.
+Second, and more importantly, we apply ordinal quantile analysis to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent–a question that cannot be answered using conditional mean models (e.g., ordinal probit).
+We note that this is the first application of ordinal quantile modeling in the EV adoption literature, enabling us to capture heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase).
+The results from ordinal probit models indicate that greater exposure to EV-related information, stronger perceptions of EVs’ environmental benefits, and higher confidence in future charging infrastructure development are positively associated with EV purchase intent, whereas perceptions of stronger government climate action are negatively associated with intent across samples.
+These relationships persist after controlling for standard demographic characteristics: younger, urban, more educated, and higher-income individuals are more likely to express interest in EVs, while women, US-born, and Black respondents report lower intent.
+The results from ordinal quantile analysis uncover substantial heterogeneity in covariate effects across the distribution of EV purchase intent.
+This approach provides richer insight into how different factors shape adoption intent.
 For example, it shows that confidence in infrastructure development and belief in environmental benefits are not only associated with higher intent among likely adopters, but also with a lower probability of reporting no interest among more skeptical respondents.
-Note that no classical estimation methods exist for ordinal quantile models, so we use a Bayesian approach.
-Finally, our use of a Bayesian framework provides several advantages over traditional (frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters.
-This can bias covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly reflect changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
-In contrast, Bayesian framework allows us to obtain posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables making direct probability statements about the parameters-rather than relying on the notion of statistical confidence as in the classical framework.
+Finally, we identify notable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
+We conclude this section by motivating our choice of a Bayesian approach for estimating the ordinal models.
+In non-linear models such as the ordinal probit, regression coefficients do not directly correspond to covariate effects–defined as changes in outcome probabilities–and these effects must therefore be computed separately for each outcome category.
+Proper estimation of covariate effects requires accounting for uncertainty in both observed variables and model parameters.
+But the latter is not captured in the Classical framework resulting in bias in covariate effects [27].
+In contrast, the Bayesian framework provides a full posterior distribution over the parameters and allows uncertainty to be systematically incorporated when computing covariate effects.
+Moreover, our adoption of Bayesian methods is driven by methodological necessity: to the best of our knowledge, frequentist estimation procedures for ordinal quantile models are not currently available, making Bayesian estimation essential in this context.
 The remainder of the article is structured as follows.
@@
 As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity.
-Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023).
+Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as [8].
 Third, the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase.
@@
 3.1 The Ordinal Probit Model
-Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates.
-These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al.
-(2023).
+Ordinal data models [30], also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates.
+These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman [26] and Batham et al.
+[5].
 In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked, without any cardinal interpretation.
@@
 Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
-The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
+The ordinal regression model, following Albert and Chib [2], can be written in terms of a continuous latent variable $z_{i}$ as follows:
 $z_{i}=x^{\prime}_{i}\beta+\varepsilon_{i},\hskip 54.2025pt\forall\;i=1,\cdots,n,$ (1)
@@
 As shown, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so forth.
-To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively (Jeliazkov et al., 2008; Jeliazkov and Rahman, 2012).
+To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively [25, 26].
 […]
@@
 But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling.
-So, following Albert and Chib (2001), we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
+So, following Albert and Chib [3], we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
 To complete the Bayesian framework, we utilize the Bayes theorem to combine the complete data likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$.
 The resulting complete data posterior can be written as,
-$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
-where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors.
+$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f_{N}(z|X\beta,I_{n})\pi(\beta)\pi(\delta),\end{split}$ (4)
+where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors.
 Given the complete data posterior (4), the conditional posteriors can be derived and the model estimated using MCMC sampling as outlined in Algorithm 1.
@@
 We evaluate the distribution of the difference $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$} for $1\leq j\leq J$, by marginalizing over $x_{i,-l}$} and $(\beta,\delta)$.
-This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
+This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects [27].
 Algorithm 1: MCMC algorithm for estimating ordinal probit model.
@@
 […]
-To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition (Jeliazkov et al., 2008).
+To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition [25].
 This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference $\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)$}, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$.
@@
 3.2 The Ordinal Quantile Model
-The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates.
-However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions.
+The parameter estimates from the ordinal probit model characterize the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates.
+However, policy-relevant questions often require examining how covariates affect specific parts of the outcome distribution, particularly the lower and upper extremes.
 For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs.
 However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers.
-To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression.
-The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
+To capture such heterogeneity in covariate effects across different levels of purchase intent, we employ ordinal quantile regression.
+The ordinal quantile regression model, proposed by Rahman [46], can be written in terms of a latent variable $z_{i}$ as follows:
 $z_{i}$ $=x^{\prime}_{i}\beta_{p}+\varepsilon_{i},$ $\qquad\forall\;i=1,\cdots,n,$ (6)
 $\gamma_{p,j-1}$ $<z_{i}\leq\gamma_{p,j}\;\longleftrightarrow\;\emph{$y_{i}$ = j},$ $\qquad\forall\;i=1,\cdots,n;\;j=1,\cdots,J,$
-where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005).
+where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution [57].
 The AL distribution is essential to create a working likelihood function, and yields the following expression,
-$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg]^{I(y_{i}=j)},\end{split}$ (7)
+$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg]^{I(y_{i}=j)},\end{split}$ (7)
 where $F_{AL}(\cdot)$ denotes the cdf of an AL distribution and $I(y_{i}=j)$ is an indicator function.
@@
 Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors.
-So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$.
+So, following Kozumi and Kobayashi [32], we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$.
 The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$.
@@
 By Bayes theorem, we combine the complete data likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at,
-$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big\{\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big\}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg\{\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg\}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
+$\begin{split}&\pi(z,\beta_{p},\delta_{p},w|y)\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big\{\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big\}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg\{\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg\}\pi(\beta_{p})\pi(\delta_{p}),\end{split}$ (8)
+where $\pi(\beta_{p})$ and $\pi(\delta_{p})$ denote the normal priors.
 In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$.
@@
     $\alpha_{MH}(\delta_{p},\delta^{\prime}_{p})=\min\bigg\{1,\frac{\;f(y|\beta_{p},\delta^{\prime}_{p})\;\pi(\beta_{p},\delta^{\prime}_{p})}{f(y|\beta_{p},\delta_{p})\;\pi(\beta_{p},\delta_{p})}\bigg\};$
-    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
+    else repeat value $\delta_{p}$. The variance of $s$ may be tuned to achieve desired acceptance rate.
   4. 4.
@@
 The conditional posterior distributions are derived from the complete data posterior (i.e., equation 8) and the parameters are sampled as per Algorithm 2.
-The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
-After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in (Rahman and Vossmeyer, 2019) and Bresson et al.
-(2021).
+The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman [41].
+After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in [45] and Bresson et al.
+[7].
 Assuming a setup similar to Section 3.1, our goal is to draw from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$.
@@
 These priors are relatively diffuse, allowing the data to primarily drive the posterior inference.
-Table 2 also reports the logarithm of marginal likelihood, which is the most accepted quantity for Bayesian model comparison (Chib and Jeliazkov, 2001; Jeliazkov et al., 2008).
-The marginal likelihood (also known as model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution.
-It accounts for both model fit and complexity, with higher values indicating a better model fit.
-The use of marginal likelihood contrasts with those commonly used in the classical framework, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) (Lin and Tan, 2017; Kumar et al., 2025).
+Table 2 also reports the logarithm of marginal likelihood [13, 25], which can be used for Bayesian model comparison.
+The marginal likelihood (or model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution.
+It accounts for both model fit and complexity, with higher values indicating stronger support for a model [^6] In Table 2, we fit an ordinal probit model (say M1) to the 2021 survey data and obtain logarithm of marginal likelihood as $\ln f(y|M1)=-12415$.
+Suppose, we now estimate a reduced model (denoted M0) that excludes the regional indicators, for which the logarithm of marginal likeilhood is $\ln f(y|M0)=-12425$.
+Then the logarithm of Bayes Factor (BF) in favor of M1 relative to M0 is, $\ln(BF)_{10}=-12415-(-12425)=10$.
+This constitutes strong evidence in favor of the model with regional indicators.
+We note that marginal likelihoods (and hence Bayes factors) are only comparable across models fitted to the same dataset; consequently, comparisons across different survey years are not meaningful, as the models are conditioned on different observed data..
+The use of marginal likelihood contrasts with commonly used classical measures, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) [38, 34].
 In ordinal models, coefficient signs indicate the direction of covariate effects for the first and last response categories.
@@
 These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories.
-Our findings complements Zhao et al.
-(2022) and Scherrer (2023), who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al.
-(2013), who underscore the role of accurate information in shaping perceptions.
+Our findings complements [58] and Scherrer [49], who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al.
+[33], who underscore the role of accurate information in shaping perceptions.
 Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability.
@@
 This finding complements Carley et al.
-(2013), Krause et al.
-(2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs.
-Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^7] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
+[10], Krause et al.
+[33], Scherrer [49], and Mamkhezri [42] who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs.
+Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^8] Relatedly, Buhmann and Criado [8] find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
 Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood ($\ln$ ML), from April 2021; May 2022; and May–June 2023 survey data.
@@
 This finding complements research, such as Carley et al.
-(2013) and Li et al.
-(2017), which highlight the importance of existing infrastructure.
+[10] and Li et al.
+[37], which highlight the importance of existing infrastructure.
 Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest.
@@
 This pattern, also noted by Carley et al.
-(2013) and Buhmann and Criado (2023), suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs.
+[10] and Buhmann and Criado [8], suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs.
 However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in EVs.
@@
 Our finding aligns with Zhao et al.
-(2022), Carley et al.
-(2013), Krause et al.
-(2013), Scherrer (2023), and Mamkhezri (2025).
+[58], Carley et al.
+[10], Krause et al.
+[33], Scherrer [49], and Mamkhezri [42].
 For instance, Zhao et al.
-(2022) finds that EV buyers in Shanghai are mostly aged 31-40.
+[58] finds that EV buyers in Shanghai are mostly aged 31-40.
 Results from the 2022 and 2023 surveys also show that female respondents are are 2.3–2.4 percentage points less likely to express strong interest, and 2.7–3.0 percentage points more likely to report no interest, than males.
 This finding is consistent with Carley et al.
-(2013), Krause et al.
-(2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men show stronger preference for EVs.
+[10], Krause et al.
+[33], Scherrer [49], Buhmann and Criado [8], and Mamkhezri [42], which also find that men show stronger preference for EVs.
 Higher education is also linked to greater EV interest.
@@
 Such a relationship is similarly documented in Carley et al.
-(2013), Krause et al.
-(2013), Buhmann and Criado (2023), and Mamkhezri (2025).
+[10], Krause et al.
+[33], Buhmann and Criado [8], and Mamkhezri [42].
 Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest.
-This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report higher EV preference among urban residents.
+This finding aligns with Buhmann and Criado [8] and Mamkhezri [42], who also report higher EV preference among urban residents.
 Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest than those in the Northeast; those in the Midwest and South show no significant difference from Northeastern respondents.
@@
 Other non-white respondents show similar preferences to White respondents.
-These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access (Hsu and Fingerman, 2021).
+These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access [21].
 Political affiliation is a consistent and significant predictor of EV purchase intent.
-Democrats are 13.9–18.8 percentage points more likely to report strong interest in purchasing EVs and 12.1–18.8 percentage points less likely to report no interest than Republicans.
-Independents and others are about 8 percentage points more likely to express strong interest than Republicans.
-These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible, as documented by Sintov et al.
-(2020).
+Relative to Republicans, Democrats are 13.9–15.6 percentage points more likely to report strong interest in purchasing an EV and 12.1–18.8 percentage points less likely to report no interest, while Independents and others are about 8 percentage points more likely to express strong interest.
+These findings align with Mamkhezri [42] and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible [53].
 They also reflect the stark policy differences between the two major political parties, particularly during the tenures of Presidents Biden and Trump.
@@
 Our results find support in Zhao et al.
-(2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals.
+[58] and Mamkhezri [42], who report greater support for EVs among higher-income individuals.
 The small covariate effect may reflect the growing availability of EVs across various market segments and the subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
@@
 4.2 Results from Ordinal Quantile Models
-We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates.
+We now present the results from ordinal quantile models [46] that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates.
 To our knowledge, this aspect has not been examined in prior work on EVs.
-Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison (Maheshwari and Rahman, 2023).
+Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison [41].
 Note that the coefficients in Table 4 are not directly comparable to those in Table 2 for two main reasons: (a) the models operate on different scales due to differences in the variance of the error terms, and (b) they address distinct inferential goals-the ordinal probit model focuses on the average probability of outcomes, whereas the ordinal quantile model targets the probability of outcomes at specific quantiles.
@@
 5 Discussion
-Our findings demonstrate how targeted strategies may be able to accelerate EV adoption.
-Exposure to EV information is associated with a 10.2 percentage point higher probability of reporting strong intent, with larger differences among respondents at the upper end of the intent distribution.
-It is also associated with an 8.5 point lower probability of reporting no intent.
-However, only 37.7% of respondents report high exposure, pointing to an opportunity for broader outreach through media, industry, and education.
-The smaller covariate effect of prior exposure at the 20th percentile suggests that simply raising awareness may not persuade consumers with low purchase interest.
-In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may influence purchase decisions.
-Belief that EVs are better for the environment than gas-powered vehicles is associated with a 15.5 percentage point higher probability of reporting strong intent, with larger differences at the upper end of the distribution.
-It is also associated with a 19.5 point lower probability of reporting no interest, with larger differences at the lower end of the distribution.
-This pattern suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters and reduces skepticism among consumers with lower purchase interest.
-In the 2021 survey, 71.8% of respondents expressed this belief.
-Emphasizing environmental benefits—such as improved air quality—may resonate with both likely adopters and more skeptical respondents, broadening support for EV adoption.
-Confidence in charging infrastructure is one of the strongest predictors of purchase intent.
-Respondents who express confidence that the US will build sufficient infrastructure are 19.1 percentage points more likely to report strong intent and 18.8 points less likely to report no interest.
-Quantile estimates suggest that confidence both strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent.
-Despite the importance of this factor, only 17.5% of respondents in the 2023 survey reported high confidence.
-Policymakers and industry can address this gap through continued investment in charging infrastructure-such as the National Electric Vehicle Infrastructure (NEVI) program-to improve affordability, accessibility, and reliability, as well as through effective communication of these efforts and of a long-term vision for the charging network.
-Belief that the government is doing too much to address climate change is associated with a 14.8–19.7 point lower probability of strong intent and a 19.7–33.4 point higher probability of no interest.
-In the 2023 survey, 24.1% of respondents held this view.
-These findings suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonate across ideological lines.
-Rather than positioning EVs primarily as a climate solution, highlighting other benefits—such as long-term cost savings and air quality improvements—may engage a broader range of potential buyers.
-Our finding that Democrats are more likely to express stronger interest in EVs, further highlights the role of political identity and the value of emphasizing non-partisan benefits to broaden appeal.
-Moreover, for consumers wary of climate-focused policies, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns.
-The results for other covariates point to further implications.
-Prior EV ownership is the strongest predictor of purchase intent, underscoring the importance of reducing barriers for first-time buyers.
-Potential options include subsidies, test-drive programs, and community-based education.
-Younger, urban, educated, and higher-income individuals report greater interest in EVs, suggesting that targeting information or advertising to such individuals may help raise adoption rates.
-By contrast, lower interest among women, US-born, and Black respondents may reflect cultural, structural, or informational barriers that could be addressed through inclusive policy design and outreach.
+Our findings point to several targeted strategies that may accelerate EV adoption by clarifying how beliefs and expectations shape purchase intent.
+First, exposure to EV-related information is associated with approximately a 10 percentage point higher purchase intent.
+This pattern is notable given that fewer than 40 percent of respondents report high exposure, suggesting scope to expand the reach of EV-related information among prospective buyers.
+At the same time, smaller estimated effects at the lower end of the intent distribution indicate that increased awareness alone is unlikely to materially increase purchase intent among consumers with low baseline interest.
+Together, these results suggest that information provision may be most effective when targeted toward consumers already considering an EV.
+Such information could focus on practical considerations such as costs, incentives, and charging logistics.
+These results complement prior work that infers information exposure through peers or media [49, 58] by relying on respondents’ self-reported exposure and by highlighting heterogeneity across the intent distribution rather than average effects.
+Second, belief that EVs have lower environmental impacts than gasoline vehicles is associated with higher purchase intent and lower reported disinterest.
+This result complements prior work linking pro-environmental attitudes to EV support [e.g., 22, 42] by showing that adoption decisions are shaped not only by general environmental orientation, but also by consumers’ perceptions of EVs’ environmental benefits relative to gasoline vehicles.
+It also extends evidence from smaller-sample studies examining the effects of perceived environmental benefits relative to gasoline vehicles to a larger, nationally representative US sample [36, 29].
+Survey responses indicate that a large majority of respondents believe EVs are environmentally preferable, but do not capture beliefs about the magnitude of those benefits.
+This measurement limitation prevents us from assessing how much additional adoption could be encouraged by shifting beliefs, and highlights a need for future research on how consumers perceive the size of environmental benefits and how those perceptions respond to information.
+Third, confidence that sufficient charging infrastructure will be built is strongly associated with purchase intent, yet fewer than 20 percent of respondents report high confidence.
+Against the backdrop of prior work emphasizing physical access to charging infrastructure [10, 55], our results point to expectations about the future charging network as a distinct channel shaping adoption decisions.
+These findings underscore the role of policy stability and credible long-term commitments in shaping expectations about the development of the charging network.
+Concerns about charger reliability and maintenance may also weaken confidence in future access even as new infrastructure is deployed.
+Such concerns were salient during the period covered by our survey data, and industry surveys reported that a substantial share of public-charging visits did not result in a successful charge due to charger unavailability or inoperability [24].
+Policymakers and industry can help address this confidence gap through continued investment in charging infrastructure—such as through the National Electric Vehicle Infrastructure (NEVI) program—together with clear communication of deployment and maintenance plans.
+Fourth, belief that the government is doing too much to address climate change is associated with lower EV purchase intent and is held by a non-trivial share of respondents.
+Our results complement prior findings linking EV interest to political affiliation [42, 53] by indicating that evaluations of climate policy itself, in addition to partisan affiliation, shape EV adoption decisions.
+This pattern implies that framing EVs primarily as a climate policy instrument is unlikely to resonate with consumers who are skeptical of climate-focused policies.
+For these consumers, information about other EV attributes—such as operating costs, performance, or local air quality—that may be less salient ex ante can still influence adoption decisions.
+Moreover, information conveyed by non-governmental actors may carry greater credibility than government-led messaging.
+Finally, prior EV ownership is strongly associated with stated purchase intent, consistent with learning or reduced uncertainty following initial adoption.
+This pattern suggests that policies lowering barriers for first-time buyers may influence adoption dynamics beyond the initial purchase.
+At the same time, systematic differences in stated interest across demographic groups may point to disparities in exposure, access, or information, highlighting the importance of inclusive policy design and outreach.
 6 Conclusion
-EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts (Bhat et al., 2025).
-However, federal support has varied along partisan lines, with Democratic administration promoting EV adoption and Republican administration prioritizing deregulation and rolling back environmental standards.
-Amid these policy shifts, consumer interest remains a key driver of both market trends and national policy.
-Although public opinion is known to influence US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of covariates.
-In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
-This study addresses these gaps by analyzing public attitudes toward EVs using three nationally representative Pew Research Center surveys from 2021-2023.
-We take the Bayesian route and model EV purchase intent with ordinal probit and ordinal quantile regression, incorporating novel predictors, prior EV ownership, political affiliation, and demographic factors.
-Results from the ordinal probit models show that strong interest in purchasing an EV is associated with greater exposure to EV information, belief in their environmental benefits, prior EV or hybrid ownership, and confidence in charging infrastructure.
-Additional positive predictors include being under 50, higher education and income, metropolitan residence, living in the West, and Democratic Party affiliation.
-Conversely, lower interest is linked to skepticism about government climate efforts, identifying as female or Black, living in the Midwest, and US-born status.
-Ordinal quantile analysis reveals that the effects of key factors on EV purchase intent vary across the distribution.
-Respondents who are knowledgeable about EVs, view them as environmentally beneficial, have confidence in charging infrastructure, own an EV, or identify as Democrats are more likely to express strong interest at higher quantiles of EV purchase intent.
-These same factors are associated with lower likelihood of no interest at lower quantiles.
-Conversely, those who believe the government is overreaching on climate policy are more likely to express no interest at lower quantiles and less likely to show strong interest at higher ones.
-These findings highlight how covariate effects differ across levels of purchase intent, thereby providing a more nuanced understanding than the ordinal probit model.
-The findings offer valuable guidance for policymakers, industry stakeholders, and advertisers seeking to bolster EV adoption.
-Confidence in charging infrastructure is one of the strongest predictors of purchase intent—associated with greater interest among likely adopters and reduced rejection among skeptics—yet less than 20% respondents report high confidence.
-Infrastructure investment, along with clearer communication about availability, reliability and expansion plans, may help close this gap.
-Exposure to EV information is also linked to stronger intent and reduced disinterest, though only 37.7% of respondents report high exposure—highlighting the value of broader outreach through media, industry, and education.
-Belief that EVs are environmentally superior to gas-powered vehicles is similarly associated with higher intent and lower disinterest, with effects varying across the intent distribution; messaging focused on benefits such as improved air quality may therefore be effective.
-Perceptions that the government is doing too much on climate change are negatively associated with intent, suggesting that outreach emphasizing cost, convenience, or energy security—particularly from non-government sources—may have broader appeal.
+Electric vehicles (EVs) are widely viewed as a cornerstone of transportation decarbonization and climate change mitigation, yet federal support for EV adoption in the US has fluctuated sharply across administrations, reflecting broader partisan divides.
+In this context, consumer purchase sentiment plays a critical role in shaping both market outcomes and the political feasibility of sustained EV policy.
+Although public opinion is known to influence US policy making, existing research has not examined EV purchase intent using nationally representative data over multiple years, nor has it quantified how certain determinants operate across different levels of purchase interest.
+This study addresses these gaps by analyzing three nationally representative Pew Research Center surveys from 2021-2023 using Bayesian ordinal probit and ordinal quantile regression models.
+The empirical results indicate that EV purchase intent is positively associated with greater exposure to EV-related information, stronger beliefs in the environmental benefits of EVs, and confidence in the development of charging infrastructure.
+Political affiliation and demographic characteristics also play a significant role, with higher intent observed among younger, more educated, and higher-income individuals, metropolitan residents, Democrats, and those living in the western United States.
+In contrast, lower purchase intent is associated with skepticism toward government climate action, as well as with certain demographic characteristics, underscoring persistent heterogeneity in attitudes toward EV adoption across population subgroups.
+The ordinal quantile analysis further reveals substantial heterogeneity in the determinants of EV purchase intent across the distribution.
+Greater EV knowledge, perceptions of environmental benefits, confidence in charging infrastructure, prior EV ownership, and Democratic Party affiliation increase the likelihood of strong purchase interest at higher quantiles while simultaneously reducing outright rejection at lower quantiles, whereas perceptions that the government is overreaching on climate policy are associated with lower interest throughout the distribution.
+These results provide a more nuanced perspective than mean-based ordinal models and yield clear policy and market insights: confidence in charging infrastructure emerges as one of the strongest predictors of adoption despite being reported by fewer than 20% of respondents, underscoring the importance of infrastructure investment and transparent communication about availability and reliability.
+Similarly, broader exposure to EV-related information and messaging emphasizing environmental benefits—such as improved air quality—are associated with higher intent and reduced disinterest, while outreach that stresses cost, convenience, or energy security, particularly from non-government sources, may be more effective among climate-policy skeptics.